### v2

# Addressing Hallucinations with RAG and NMISS in Italian Healthcare LLM Chatbots

Abstract
I combine detection and mitigation techniques to addresses hallucinations in Large Language Models (LLMs). Mitigation is achieved in a question-answering Retrieval-Augmented Generation (RAG) framework while detection is obtained by introducing the Negative Missing Information Scoring System (NMISS), which accounts for contextual relevance in responses. While RAG mitigates hallucinations by grounding answers in external data, NMISS refines the evaluation by identifying cases where traditional metrics incorrectly flag contextually accurate responses as hallucinations. I use Italian health news articles as context to evaluate LLM performance. Results show that Gemma2 and GPT-4 outperform the other models, with GPT-4 producing answers closely aligned with reference responses. Mid-tier models, such as Llama2, Llama3, and Mistral benefit significantly from NMISS, highlighting their ability to provide richer contextual information. This combined approach offers new insights into the reduction and more accurate assessment of hallucinations in LLMs, with applications in real-world healthcare tasks and other domains.
Keywords: Large Language Models, Retrieval-Augmented Generation, Hallucinations, Healthcare, Natural Language Processing, Evaluation Metrics, Information Systems
1 Introduction
Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by Vaswani [52]. Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the nuances of language. These models perform various tasks, from basic text classification to advanced text generation, and more recent iterations can even interpret images and generate code.
LLMs have surged in popularity due to their intuitive, prompt-based interfaces, enabling effortless interaction. The past few years have seen the continuous development of notable LLMs like ChatGPT [33, 1], Gemini [11], Falcon [37], Llama [49, 50, 30], and Gemma [46, 47]. As highlighted by [15], these models excel in language comprehension [14, 16], text summarization [59], and reasoning [54, 22, 38]. As a result, their applications span various industries, such as processing electronic medical records, matching clinical trials, and drug discovery in healthcare. In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [35].
However, LLMs face a significant challenge: hallucinations [12, 51]. Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [17]. This behavior is a serious threat since it affects the application of LLMs to real-world cases [20], as hallucinations can undermine the reliability of their outputs. [15] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context. Factuality hallucinations manifest as either factual inconsistencies or fabrications. For example, a model might mistakenly claim that “The Colosseum is located in Berlin, Germany” instead of Rome or invent details about non-existent technologies. Faithfulness hallucinations occur when the output deviates from user instructions or context. These can involve instruction inconsistencies (e.g., responding in Spanish when asked to translate into Italian), context inconsistencies (introducing unrelated information), or logical errors (e.g., miscalculating “two times four” as three).
The mitigation of hallucinations has become an urgent concern due to the widespread use of LLMs in sensitive areas such as healthcare. Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [34], their opaque black-box nature that hides knowledge boundaries [45, 42], and their training on large datasets that may contain outdated or incorrect information [60]. Additionally, their versatility across many tasks complicates mitigation efforts, while their strong performance can make hallucinations, especially factual fabrications, harder to detect.
Researchers focused on various techniques to mitigate hallucinations. One of the most promising methods is Retrieval-Augmented Generation (RAG) [24], which incorporates external knowledge during text generation to anchor outputs in factual information. RAG addresses issues of accuracy and currency in LLM responses by using authoritative sources, generating verifiable answers, and enhancing user confidence across applications [21]. Other approaches include Prompt Engineering that involves optimizing LLMs outputs with different instructions [48, 55], self-refinement through feedback and reasoning [27], where feedback on LLM outputs improves the accuracy and quality of subsequent responses, and prompt tuning, which modifies instructions during fine-tuning to improve task-specific performance [23].
Mitigating hallucinations requires first their detection, and various metrics have emerged. Traditional metrics like ROUGE [26], BLEU [36], and METEOR [4] measure consistency between generated text and source text, but they often struggle with semantically relevant but superficially divergent text [5]. To address these limitations, newer embedding-based methods such as BERTScore [58] use contextual embeddings to assess semantic similarity between sentences. Another approach involves ”opening the box” where token probabilities or logit outputs of LLMs are analyzed to detect hallucinations [3, 51]. However, these techniques require access to internal model states, which is not always feasible for proprietary systems.
This paper proposes a general framework for detecting and mitigating hallucinations, with an application to the Italian healthcare domain. For the mitigation part, I employ a RAG architecture to integrate external knowledge into the generation process, ensuring that the models are grounded in factual data. The methodology applies a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. I compare a range of LLMs, including Gemma [46], Gemma 2 [47], Mistral [18], GPT-4 [1], Llama2 [50], and Llama3 [30] for the answer generation. This framework uses a prompting strategy to instruct models to generate answers based on the provided context. To assess interpretability and difficulty, I categorize 100 questions into three levels of understanding: low, mid, and high. For the detection strategy, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring framework designed to be applied alongside existing metrics to integrate contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS refines traditional metrics by highlighting cases where they may unfairly penalize performance despite the enhanced contextual relevance of the responses.
Results show that Gemma2 and GPT-4 achieve the highest overall performance in Italian healthcare-related question-answering tasks across all difficulty levels, while Gemma exhibits the highest rate of hallucinations. The NMISS scoring system provides a more accurate assessment of LLM performance, especially in high-difficulty tasks for mid-tier models like Llama2, Llama3, and Mistral, which show notable improvements compared to traditional metrics. Additionally, NMISS recognizes the value of contextual information provided by models in cases where they fail to deliver exact matches to expected answers in difficult tasks. By focusing on the relevance and coherence of the provided context, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they do not fully align with the reference answers.
This paper is organized as follows: Section 2 reviews the relevant literature on hallucination detection and mitigation. Section 3 details the methodology, highlighting the contributions of NMISS. Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 presents my findings. Finally, Section 6 summarizes contributions and outlines future research directions.
2 Related Works
2.1 Hallucination Detection Techniques
Hallucinations in LLMs have been widely studied. [15] provides a detailed overview of strategies to detect and mitigate hallucinations, offering insight into their occurrence in current LLMs architectures. Ensuring LLM’s trustworthiness depends on early detection, as emphasized by [51], who highlights the importance of proactive measures during text generation.
[60] categorize detection methods into three main types. First, logit-based estimation, which uses token-level probabilities or entropy from the model’s logits [39]. Second, verbalized confidence estimation, where LLMs are asked to self-rate the confidence in their responses. Third, consistency-based estimation involves generating multiple candidate outputs and comparing them to detect logical inconsistencies, mimicking human reasoning to refine answers [53, 44, 61].
Furthermore, [19] categorize hallucination detection into white-box, grey-box, and black-box techniques. White-box methods analyze the model’s internal workings to identify when incorrect statements are generated. Grey-box approaches examine logit outputs to find areas of low confidence in the output. Black-box techniques, which only require the model’s final answers, are simpler to apply since they don’t require access to internal states.
Recent studies introduce SAPLMA [3], which detects hallucinations by using a classifier trained to predict the truthfulness of statements based on LLMs’ hidden layer activations. This approach outperforms traditional methods that rely solely on sentence probabilities. [51] propose an iterative process that detects potential hallucinations by identifying tokens with low normalized probabilities. The system then retrieves relevant information to correct these errors, effectively reducing hallucinations in GPT-3.5.
2.2 Resource-Intensive Mitigation Strategies
In addition to detection, various strategies have been developed to mitigate hallucinations in LLMs. One common approach is Knowledge Augmentation, such as Knowledge Injection (KI). [29] enhance prompts with information from knowledge graphs, such as adding details like location names for review tasks. This significantly reduces hallucinations and improves factual accuracy, with KI-LLMs producing more correct assertions and outperforming larger models in response quality.
[31] take a direct approach by training T5 models on knowledge graph triples from Wikidata. By masking key entities, the model learns and generalizes knowledge effectively. SKILL, after fine-tuning, outperforms standard T5 models on question-answering tasks, proving the effectiveness of structured data infusion.
Another key mitigation strategy is retrieval-based frameworks. [25] combines a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework. This integration enhances the accuracy and reliability of generated content. Similarly, [13] apply retrieval techniques in biomedical contexts, incorporating sources like UMLS and Wikipedia to generate more factual summaries, which are crucial for healthcare applications.
[62] introduce a self-supervision framework that assigns risk scores, called POLAR scores, to LLM responses as indicators of hallucination risk. Their approach uses Pareto optimization to balance supervision from multiple sources. The process involves gathering LLM responses, training a probabilistic function $h$ to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support. Experimental results show significantly reduced hallucination rates across tasks like question answering and text generation.
[57] propose the MixAlign framework, which aligns user queries with external knowledge to resolve ambiguities. This improves response accuracy and reduces hallucinations, with results showing a sharp increase in correct answers and a decrease in hallucinations.
2.3 Zero-Resource Hallucination Strategies
The strategies discussed in Section 2.2 are resource-intensive. In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited resources, relying only on model outputs without needing internal access or external data. Many of these techniques not only detect hallucinations but also help mitigate them by refining responses based on internal consistency assessments.
One approach is through self-consistency techniques, which compare multiple responses generated by the LLM itself. [28] propose a zero-resource method that generates multiple responses from GPT-3 and compares them to a reference answer to assess factual consistency. This method detects and mitigate hallucinations using low-level metrics and outperforms traditional grey-box and black-box methods.
Another approach is self-rating techniques, as proposed by [6]. The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers. RelD is trained by converting a regression task into a binary classification task, aligning with human intuition. By flagging low-confidence responses, it serves as an indirect mitigation mechanism. Experiments across various datasets and models (e.g., BLOOM [56]) show RelD’s effectiveness, validated by both automatic metrics and human-in-the-loop evaluations. Similarly, [2] addresses hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources. DQs involve asking the model a binary question multiple times to calculate a probability. In contrast, IQs ask about the authors or content of the reference and check for consistency across responses. Although primarily designed for detection, these methods can also reduce hallucinations by discarding low-confidence responses. Results show that IQs generally perform better than DQs in detecting hallucinations, though performance varies among models like GPT-4, GPT-3, and Llama 2.
Finally, [32] explore self-contradiction techniques as a factor in hallucinations. They introduce a prompting framework to detect and mitigate contradictions within the same response, relying solely on the model’s logical reasoning capabilities. This framework includes functions for gathering contextual information, generating sentences, identifying contradictions, and refining sentences. Across various LLMs, this approach achieves an 80% F1 score.
2.4 Metrics for Evaluating Hallucinations
Traditional metrics to detect or manage hallucinations [5] are notably ROUGE [26], BLEU [36], and METEOR [4]. These indicators measure the consistency between the generated text and the source text. However, a drawback of these simple metrics is that they do not accurately capture distance word-dependencies or if the two answers diverge in lexicon but have equal meaning. BERTScore overcomes this issue [58] by leveraging pre-trained contextual embeddings from BERT [9] and matches words in candidate and reference sentences by cosine similarity, correlating well with human judgments.
2.5 Contributions
Despite significant progress in hallucination detection and mitigation, key challenges persist. Many existing methods, such as knowledge injection or retrieval, are resource-intensive, limiting their practicality in certain applications. Additionally, traditional evaluation metrics often fail to account for contextual accuracy, unfairly penalizing semantically correct responses that differ lexically from reference answers.
This paper addresses these gaps in several steps. The key contributions of this work are as follows:
(1) A combined approach for hallucination detection and mitigation in LLMs. I develop a RAG-based architecture to integrate external knowledge into the generation process, ensuring that models are grounded in factual data. The methodology leverages a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. This framework includes a prompting strategy to instruct models to generate answers based on the provided context, enabling a comparative evaluation of various LLMs using this approach.
(2) Comprehensive analysis of hallucinations across reasoning levels. To evaluate hallucination tendencies in LLMs, I analyze their performance across progressively challenging levels of reasoning, categorized in low, mid and high levels. This approach enables an assessment of how hallucination rates and mitigation effectiveness vary as tasks demand increasing degrees of logical complexity and understanding. The progressive evaluation highlights how model reliability and factual grounding are affected under escalating reasoning challenges, providing insights into their strengths and vulnerabilities.
(3) Development of a detection scoring system. I introduce the Negative Missing Information Scoring System (NMISS), a novel framework that fine-grains traditional metrics by incorporating contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS highlights cases where traditional metrics may penalize models despite the enhanced contextual relevance of their answers.
3 Methodology
3.1 RAG Architecture
RAG models combine the strengths of retrieval-based and generative language models to address issues like hallucinations and factual inaccuracies by integrating a retrieval mechanism to ground responses in relevant information [24].
I formally define the problem as follows. Let $q$ represent the query, $c$ the context information retrieved from external sources, and $r$ the output sequence generated by the model. The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-K set of documents $k$ based on the query $q$, and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
$p_{\text{RAG}}(r\mid q)=\sum_{c\in k(p(\cdot\mid q))}p_{\eta}(c\mid q)\prod_{i=1}^{N}p_{\theta}\left(r_{i}\mid q,c,r_{1:i-1}\right)$ (1)
This RAG-sequence approach ensures that the response remains consistent by using the same retrieved document to generate each token, maintaining coherence throughout the sequence. Here, $N$ represents the length of the output sequence. This architecture is particularly advantageous for domains like healthcare, where factual accuracy is critical. By leveraging retrieval from medical documents, guidelines, or research articles, the RAG model ensures that responses are accurate, minimizing the risks associated with misinformation [10]. In this paper, I employ the retrieval and generation models separately and combine them for inference; the workflow is detailed in Section 4.2.
3.2 Evaluation Metrics
The metrics described in this section are used to evaluate the quality of responses generated by the system, referred to as the System Summary (SS), against the ground truth responses, referred to as the Reference Summary (RS), which are human-generated responses in this study. Both SS and RS are treated as sets of n-grams for all metrics. Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$, where $s_{i}$ is an n-gram in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$, where $r_{j}$ is an n-gram in the RS.
ROUGE [26] evaluates responses based on precision, recall, and F1-score. The text is pre-processed before calculating these scores [43]. More formally, ROUGE-N measures the overlap of n-grams between the SS and RS. The precision, recall, and F1-score for ROUGE-N are defined as follows:
$\text{Precision}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{SS}}|}$ (2)
$\text{Recall}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{RS}}|}$ (3)
[…]
where $S_{\text{SS}}$ is the set of n-grams in the SS, $S_{\text{RS}}$ is the set of n-grams in the RS, $|S_{\text{SS}}\cap S_{\text{RS}}|$ is the cardinality of the intersection of $S_{\text{SS}}$ and $S_{\text{RS}}$, $|S_{\text{SS}}|$ is the cardinality of the set $S_{\text{SS}}$, and $|S_{\text{RS}}|$ is the cardinality of the set $S_{\text{RS}}$. Precision measures the accuracy of the generated text, while recall evaluates the coverage of the information. Low precision suggests hallucination, while low recall indicates a coverage issue.
Another variant used here is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS. The LCS represents the longest subsequence in both SS and RS while preserving word order, though the words do not need to appear consecutively [7].
BLEU [36] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers. BLEU accounts for variability in word choice and word order found in different, yet valid, translations of the same source sentence. It quantifies the overlap by measuring how many n-grams from the SS are present in the RS. The core of the BLEU metric is the modified n-gram precision, which addresses the over-generation problem by treating each reference word as exhausted once a matching candidate word is found.
Let $V$ be a function that counts the occurrences of an n-gram in a given set of n-grams. Formally, for an n-gram $w$ and a set of n-grams $S$, the function $V$ is defined as $V(w,S)=|\{s\in S\mid s=w\}|$. This function returns the number of times $w$ appears in the set $S$. The modified n-gram precision $p_{n}$ is calculated as:
$p_{n}=\frac{\sum_{w\in S_{\text{SS}}}\min\left(V(w,S_{\text{SS}}),\max_{j}(V(w,S_{\text{RS}_{j}}))\right)}{\sum_{w\in S_{\text{SS}}}V(w,S_{\text{SS}})}$ (5)
where $V(w,S_{\text{SS}})$ is the number of times the n-gram $w$ appears in SS, and $\max_{j}(V(w,S_{\text{RS}_{j}}))$ is the maximum number of times $w$ appears across all RS. Here, $j$ iterates over the different reference summaries in $S_{\text{RS}}$, where $j=1,2,\dots,m$, and $m$ is the number of RS.
To account for translations that are too short, BLEU applies a brevity penalty (BP), defined as:
$\text{BP}=\begin{cases}1&\text{if }b>r\\ e^{(1-r/b)}&\text{if }b\leq r\end{cases}$ (6)
where $b$ is the length of the candidate translation (SS), and $r$ is the effective reference length (RS).
The final BLEU score is computed by taking the geometric mean of the modified n-gram precisions (up to length $N$, typically set to 4), and multiplying by the brevity penalty:
$\text{BLEU}=\text{BP}\cdot\exp\left(\frac{1}{N}\sum_{n=1}^{N}\log p_{n}\right)$ (7)
METEOR [4] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms. METEOR computes the score by comparing the SS to each RS and reporting the highest score. Additionally, METEOR introduces a Discontiguous Matches Penalty (DP), based on the number of matched chunks.
Let $T$ represent the number of contiguous chunks of matched unigrams between SS and RS. For a set of matched unigrams $V=\{v_{1},v_{2},\dots,v_{k}\}$, a chunk is defined as a maximal sequence of consecutive matches $v_{i},v_{i+1},\dots,v_{i+t}$, such that $v_{i+t}=v_{i}+t$ for all $t\geq 0$.
The function $T(V,S_{\text{SS}})$, which counts the number of contiguous chunks, is then formalized as:
$T(V,S_{\text{SS}})=\left|\{\{v_{i},v_{i+1},\dots,v_{i+t}\}\subseteq V\mid v_{i+t}=v_{i}+t\}\right|$ (8)
This function returns the number of contiguous chunks of matched unigrams in $S_{\text{SS}}$. The Discontiguous Matches Penalty (DP) is then computed as:
$\text{DP}=0.5\cdot\left(\frac{T(V,S_{\text{SS}})}{|V|}\right)$ (9)
where $|V|$ is the number of matched unigrams. The penalty increases as the number of chunks grows, penalizing fragmented matches. The final METEOR score is adjusted by applying DP to the $F_{\text{mean}}$ score:
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{DP})$ (10)
BERTSCORE is an evaluation metric for text generation that leverages pre-trained BERT embeddings [58]. Unlike traditional metrics that rely on exact n-gram matches, BERTSCORE measures the semantic similarity between tokens in the System Summary (SS) and Reference Summary (RS) by comparing their contextual embeddings. This method captures the semantic meaning even when the exact words differ. BERTSCORE generates contextual embeddings for the tokens in both the SS and RS, where each embedding reflects the surrounding context of the word. The similarity between tokens is computed using cosine similarity. For each token in the RS, BERTSCORE identifies the most similar token in the SS, and vice versa, to compute precision and recall.
Given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token. The similarity between two embeddings is computed using cosine similarity. Precision ($P_{\text{BERT}}$) and recall ($R_{\text{BERT}}$) are calculated as follows:
$R_{\text{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}\cos(e(x_{i}),e(\hat{x}_{j}))$ (11)
$P_{\text{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}\cos(e(\hat{x}_{j}),e(x_{i}))$ (12)
The final BERTSCORE is computed as the F1 score of the precision and recall:
$F_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$ (13)
Exact Match (EM) is a straightforward metric that penalizes any response not exactly equal to the provided dataset annotation. It is analogous to conventional classification accuracy, ensuring the generated text matches the reference text precisely. The EM can be defined as:
$EM=\begin{cases}1&\text{if }SS=RS\\ 0&\text{if }SS\neq RS\end{cases}$ (14)
3.3 NMISS
3.3.1 Rationale
The metrics discussed above evaluate the quality of text summaries and machine-generated responses by comparing them to reference answers. While effective, this method overlooks the possibility that non-matching words or n-grams might still be valuable if they align with the question’s broader context. This observation prompts an investigation into whether the words or phrases that do not match the reference could still be relevant compared to the broader context in which the response is generated. LLMs often produce responses beyond the strict instructions, frequently adding extra content, as shown in Table 1. This additional content can manifest in two ways: as hallucinations, where the model generates information that is factually inaccurate or irrelevant, or as meaningful elaborations, where the model extends the response in a contextually relevant manner.
Table 1: Example of Question, Context, Reference, and System Answer
Question How to prevent influenza?
Context [1] of the Relaunch Decree to: 1. prioritize investment in prevention to minimize impacts at all levels; 2. include in the Plans for managing a possible (re)emergence of an epidemic the entirety of health and social services to create the possibility of continuity of care even in the event of a “resurgence” of the epidemic (like a “generator group”), not only in hospitals but also in the community; 3. ensure that investments, particularly in personnel, cover all measures aimed at strengthening containment measures to reduce the risk of contagion. People should internalize and adopt recommendations such as washing hands thoroughly and frequently, using masks extensively, correctly, and appropriately, avoiding unnecessary movements and behaviors linked to social and recreational activities that could create spontaneous gatherings [3] as it is essential to maintain social distancing, use masks, and sanitize hands carefully to contain contagion. It is also crucial to self-isolate in case of cold symptoms. There are no risk locations but rather risky behaviors. We must convey to citizens the awareness that if the necessary precautions are followed, it is difficult to contract and transmit this infection, and most normal social and work activities can continue. [4] For the first time, high levels of criticality regarding the virus’s spread in our country are reported. The population is urged to avoid events and gatherings in both public and private settings. Individual strict behavior is mandatory to limit transmission risks and prevent a further and faster worsening of the epidemic. The local transmission of the virus is widespread across the entire territory.
Reference I’m sorry, I can’t help you based on the information I have.
System I’m sorry, I can’t help you based on the information I have, as the question is not related to the Relaunch Decree but rather to influenza, which is a different type of illness than the epidemic under discussion. However, based on the context, it is noted that the Relaunch Decree emphasizes the importance of strict behaviors to minimize impacts at all levels, such as social distancing, mask use, and hand hygiene, which are generally useful measures for preventing disease transmission in general, not just the discussed epidemic.
By accounting for these non-matching elements, I can more effectively evaluate the performance of LLMs, distinguishing between irrelevant or hallucinated information and valuable, contextually appropriate elaborations. This approach offers a more nuanced understanding of the model’s output.
The NMISS scoring system was developed to capture the potential value in these unmatched words by evaluating whether they align with the contextual information provided. This rationale stems from the observation that LLMs often do not strictly adhere to the given instructions, incorporating additional content into their answers. This content may either be a hallucination or a contextually relevant elaboration that enriches the response.
3.3.2 Formulation
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer. Define the set of matching tokens between system-generated answers and reference answers as $\mathcal{P}(s_{i},r_{i})$, and let $\mathcal{N}(s_{i},r_{i})=s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ represent the tokens in $s_{i}$ that do not match $r_{i}$. Finally, let $\mathcal{A}(s_{i},c_{i})=\mathcal{N}(s_{i},r_{i})\cap c_{i}$, which represents the set of unmatched tokens in $s_{i}$ that match the context $c_{i}$. I aim to define a function
$f_{NMISS}:\mathcal{S}\times\mathcal{R}\times\mathcal{C}\rightarrow\mathbb{R}$ (15)
that evaluates how well system-generated answers align with both reference answers and the context. First, I define a reference-based evaluation function $f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$, and a context-based evaluation function $f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$, which computes the score based on tokens matching the context. The NMISS evaluation function is defined as:
$f_{\text{NMISS},i}(s_{i},r_{i},c_{i})=\max\left(f_{\text{ref},i}(s_{i},r_{i}),\frac{\lambda_{1}\cdot f_{\text{ref},i}(s_{i},r_{i})+\lambda_{2}\cdot f_{\text{cxt},i}(s_{i},c_{i})}{\lambda_{1}+\lambda_{2}}\right)$ (16)
where $\lambda_{1}=|\mathcal{P}(s_{i},r_{i})|$ and $\lambda_{2}=|\mathcal{A}(s_{i},c_{i})|$ represent the cardinalities of their respective sets. $\lambda_{1}$ and $\lambda_{2}$ represent the number of matching reference and context matches tokens, respectively.
To safeguard the integrity of standard scoring metrics, NMISS implements two key mechanisms: a weighted scheme and a maximum function. These mechanisms are essential for balancing reference-based and context-based contributions, addressing challenges such as score underestimation or overestimation when processing contextual information. The weighted scheme ensures a fair combination of reference-based and context-based scoring, preventing a disproportionate reliance on one over the other. Specifically, $\lambda_{1}$ and $\lambda_{2}$ control the relative influence of reference overlap and contextual alignment. This mechanism helps NMISS capture instances where a system-generated response extends beyond the reference but remains faithful to the context, distinguishing it from hallucinated content. The maximum function further enhances robustness by preventing NMISS from being unfairly penalized when context length inflates the denominator. In recall-sensitive metrics, an inflated denominator—such as the total number of n-grams in the context—could lower the score even if the additional content is meaningful. The formulation in Equation 16 ensures that context-based contributions only positively influence the score when they genuinely add value. In other words, if context matching introduces more noise than useful content, NMISS defaults to the reference-based score. This design prevents cases where large but irrelevant contexts artificially suppress the metric. This approach is particularly advantageous in tasks like question-answering, where the model’s response might not strictly match the reference but still be valid given the context. By integrating these mechanisms, NMISS ensures coherence in scoring, making it a more flexible and context-aware evaluation metric. The detailed implementation is outlined in Algorithm 1.
Algorithm 1 NMISS Algorithm
1:=$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, the set of reference answers
2:$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, the set of system-generated answers
3:$\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$, the set of context texts
4:=$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, the set of NMISS scores for a metric $\mathcal{X}$
5:
6:for each $(s_{i},r_{i},c_{i})$ in $(\mathcal{S},\mathcal{R},\mathcal{C})$ do
7:     Stage 1: Reference-Based Metric Calculation
8:     $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$triangleright Matched tokens
9:     $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$triangleright Unmatched tokens
10:     $f_{\text{ref},i}(s_{i},r_{i})\leftarrow\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$
11:
12:     Stage 2: Context-Based Metric Calculation
13:     $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$triangleright Context-matched tokens
14:     $f_{\text{cxt},i}(s_{i},c_{i})\leftarrow\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$
15:
16:     Stage 3: NMISS Calculation
17:     $f_{\text{wgt},i}(s_{i},r_{i},c_{i})\leftarrow\frac{\lambda_{1}\cdot f_{\text{ref},i}(s_{i},r_{i})+\lambda_{2}\cdot f_{\text{cxt},i}(s_{i},c_{i})}{\lambda_{1}+\lambda_{2}}$
18:     $f_{\text{NMISS},i}(s_{i},r_{i},c_{i})\leftarrow\max(f_{\text{ref},i}(s_{i},r_{i}),f_{\text{wgt},i})$
19:     $\mathcal{M}_{\text{NMISS}}\leftarrow\mathcal{M}_{\text{NMISS}}\cup f_{\text{NMISS},i}(s_{i},r_{i},c_{i})$
20:end for
4 Experimental Setup
4.1 Data
Data comprises text-based news in Italian, collected from various sources, and all focused on the health domain within Italy. This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues across the country. I collected 126,470 news articles from 2010 to 2024; their frequency is shown in Figure 1. A pre-processing step was employed to clean the raw text of HTML tags and other unwanted artifacts to ensure data quality. Token statistics are calculated to understand the overall structure of the data and are shown in Table 2. On average, each article contains approximately 348 tokens, with a standard deviation of 264 tokens, indicating some variability in article length. The shortest article consists of just 6 tokens, while the longest contains up to 5,594 tokens. The median article length is 284 tokens, with the interquartile range spanning from 195 to 421 tokens, reflecting a diverse range of article sizes.
Table 2: Summary of Token Statistics for the News Collection
N. Tokens Mean Std Min 25% Median 75% Max
44018696 348 264 6 195 284 421 5594
Figure 1: News Frequency Aggregated by Year
Next, I developed a list of 100 questions related to public health. Questions range from very generic to specific or ambiguous, leading to different levels of interpretation ease. The questions are categorized into three levels of understanding, based on the objective criterion of question complexity, measured by the average question length as a proxy for how detailed the question is (low: 30 characters, mid: 40 characters, high: 50 characters). Additionally, all questions were reviewed by a domain expert to ensure accuracy and relevance. All questions have been answered manually; Table 3 shows an example of questions.
Table 3: Categorized Example Questions
Category Example Question
HIGH How can epigenetics help us understand diseases?
MID What happens to the body when we are stressed?
LOW What is a common cold?
4.2 User-Agent Chatbot Framework
This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news, using a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing. This multi-model setup ensures robustness in handling diverse queries related to Italian health news, providing accurate, coherent, and contextually grounded responses.
The system begins with Sentence Transformers to generate dense vector embeddings for the Document Database, which are stored and accessed via the Dense Vector Space. This enables efficient retrieval of relevant information through semantic search and clustering. After experimenting with different chunk lengths, I optimized retrieval performance by selecting a chunk length of 512 tokens with a 64-token overlap, as evaluated by human reviewers on a dataset of 100 health-related queries. For the retrieval part, I choose the Sentence Transformer model ‘paraphrase-multilingual-MiniLM-L12-v2‘ [^1] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for its high-quality embeddings across multiple languages, making it ideal for Italian health news [^2] After testing various chunk token sizes, the chunk length of 512 tokens with a 64-token overlap provide the best balance between retrieval performance and contextual continuity..
Figure 2: User-Agent Application Workflow. Text data is split into chunks for efficient retrieval, indexed using a Sentence Transformer model, and stored as dense vector embeddings in a vector store. The system retrieves relevant chunks based on query embeddings and passes them to an LLM for response generation.
For the retrieval part, I choose various LLMs handling the Italian language. LLaMA2 and LLaMA3 are generative models developed by Meta[49, 30], optimized through supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF). Mistral[18] is a 7-billion parameter model offering competitive performance in generating accurate responses. GPT-4[1], the latest multimodal model from OpenAI. Finally, Gemma and Gemma 2[46, 47], lightweight models from Google, are fine-tuned for Italian question-answering and summarization tasks.
The workflow illustrated in Figure 2 starts with retrieving relevant information from a large corpus of chunk news by embedding the user query into a vector space using the sentence transformer. The query embedding is matched against stored dense vector embeddings, and the top-K most similar document chunks are retrieved from the vector database. Next, the generation component uses the retrieved document chunks as context to produce a response. The LLM receives both the original query and the retrieved context, generating a natural language response grounded in factual data, thus reducing hallucinations.
4.3 Prompt Engineering Template
I develop a question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context [^3] The list of the final set of Italian questions is available upon request.. An English version is shown in Table 4. This template ensures that the chatbot remains anchored to the information in the provided context. If the answer to a query is not found within the context, the chatbot will respond with: I’m sorry, I can’t help you based on the information I have”.
Table 4: Prompt Engineering - Example (English translation)
Prompt You are a chatbot that loves helping people! Given the following context section, answer the question using only the provided context. If you are unsure and the answer is not explicitly written in the documentation, respond by saying, ’I’m sorry, I can’t help you based on the information I have.’
Question How can myocardial infarction be prevented?
Context [1] The diagnosis of myocardial infarction on-site allows for rapid transportation by equipped ambulance, uncomfortable but not risky, directly to the hemodynamic lab, enabling coronary reopening at least 30 minutes earlier than if the patient arrives at the emergency room with their own means at the reference hospital.” – responsible for myocardial infarction [2] early infarction and saving lives with a simple blood test.” According to a press release from the Mauriziano’s press office. ”In the last ten years,” the press release explains, ”the diagnosis of acute coronary syndromes has progressively shifted attention from the electrocardiogram (always and essential) to the qualitative and quantitative biochemical definition of myocardial damage. The new method provides a more specific, sensitive, and early definition of myocardial damage with undeniable advantages in response speed to the pathology, especially when it presents subtly and unexpectedly. In short, it recognizes infarction practically in real-time.” [3] ”Remembering that acute myocardial infarction (AMI) is a time-dependent pathology,” Mura concludes, ”it is essential to prevent and initiate therapy, either pharmacological or interventional in the hemodynamic room, within the timeframes established by national protocols.
System Based on the context, there is no precise information on the symptoms indicating myocardial infarction. However, it mentions a new hospital technique capable of recognizing infarction in real-time, which is crucial given the time-dependent nature of the pathology.
5 Results
5.1 RAG Performance Analysis
The results are presented in Figures 3-4, showing the models’ performances across three levels of questioning. According to [8], a ROUGE score below 0.4 is considered low, between 0.4 and 0.5 is moderate, and above 0.5 is good.
The Gemma model shows the weakest performance across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4, indicating significant difficulty in capturing relevant unigrams and bigrams. The ROUGE-L score is similarly weak, suggesting issues with structural coherence. Additionally, Gemma’s BLEU and METEOR scores remain low, showing poor n-gram precision and recall. With no EM, Gemma struggles to generate sequences that match the reference.
In contrast, Gemma2 emerges as the top performer across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, indicating strong unigram and bigram relevance. The high ROUGE-L score points to solid structural coherence, while the BLEU and METEOR scores suggest excellent n-gram precision and recall. Gemma2 also achieves notable EM rates, ranging between 40% and 50%, highlighting its strength in generating exact matches, which is critical for tasks requiring precise language generation.
GPT-4 delivers balanced performance across all metrics but shows variability across question levels. It follows closely behind Gemma2 in ROUGE, BLEU, and METEOR scores. However, a drop occurs at mid and high question levels, suggesting GPT-4 faces challenges with complex queries. Despite this, its BERTSCORE exceeds 0.9, and it also performs strongly in EM, rivaling Gemma2 in generating exact matches.
Llama2, Llama3, and Mistral exhibit similar patterns. While their F1 scores for all ROUGE variants are relatively low, mainly due to weak precision, their recall scores approach those of Gemma2 and GPT-4. Mistral maintains consistent performance across all question levels, but the Llama models struggle more with more difficult questions.
Overall, Gemma2 emerges as the most performant model, strictly followed by GPT-4, which remains competitive despite a slight decline in performance on complex queries. Mistral and the Llama models perform modestly across ROUGE, BLEU, and METEOR, with Mistral being more stable across all levels. Gemma lags significantly behind, particularly in ROUGE and BLEU. BERTSCORE remains consistently high across all models, with Gemma2 and GPT-4 achieving BERTSCOREs above 0.9.
Figure 3: ROUGE Metrics by Question Levels
Figure 4: BLEU, METEOR and EXACT MATCH Metrics by Question Levels
5.2 NMISS Performance Evaluation
Results for BLEU and METEOR NMISS scores are shown in Tables 5-6. Findings report the scores against the novel metric for all the models across the question levels. In both cases, NMISS scores generally surpass their standard counterparts, indicating enhanced performance across different difficulty levels.
Table 5: BLEU versus BLEU NMISS scores
BLEU score BLEU NMISS score
Model LOW MID HIGH LOW MID HIGH
gpt4 0.650 0.468 0.473 0.653 0.469 0.484
llama2 0.122 0.072 0.067 0.151 0.098 0.090
llama3 0.152 0.208 0.125 0.170 0.259 0.173
mistral 0.272 0.263 0.191 0.291 0.291 0.225
gemma 0.001 0.000 0.047 0.001 0.011 0.049
gemma2 0.654 0.578 0.604 0.668 0.626 0.664
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
For both metrics, models like GPT-4 show minimal differences between standard and NMISS scores, indicating high consistency in performance across both metrics. However, models like Llama2, Llama3 and Mistral exhibit more significant improvements in their BLEU NMISS scores, suggesting that NMISS better captures the models’ ability to provide additional contextual information, especially in more challenging tasks. Notably, Gemma2 shows strong performance overall, with NMISS scores slightly higher than standard ones, particularly at mid and high levels, highlighting its ability to maintain accuracy and detail in complex scenarios. The increase in NMISS scores when models struggle to answer could be attributed to how NMISS rewards the inclusion of additional relevant information or contextual details. When a model does not provide a concise answer, it often compensates by providing surrounding context, explanations, or related information. While not a perfect match to the expected answer, this additional content contributes meaningfully to the overall response.
Table 6: METEOR versus METEOR NMISS scores
METEOR score METEOR NMISS score
Model LOW MID HIGH LOW MID HIGH
gpt4 0.714 0.526 0.565 0.714 0.526 0.567
llama2 0.443 0.325 0.291 0.453 0.330 0.310
llama3 0.415 0.442 0.336 0.429 0.455 0.355
mistral 0.431 0.485 0.451 0.436 0.488 0.456
gemma 0.010 0.012 0.073 0.011 0.019 0.080
gemma2 0.718 0.675 0.724 0.723 0.682 0.731
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
Tables 7-9 compares ROUGE and ROUGE NMISS scores across different models and difficulty levels. Similar to the trends observed with BLEU and METEOR, the ROUGE NMISS scores generally outperform the standard ROUGE metrics, especially at higher difficulty levels. This consistent pattern across different metrics suggests that NMISS captures additional relevant information, rewarding models for providing more detailed and contextually accurate responses.
Specifically, for most models, ROUGE NMISS scores surpass the standard ROUGE-1 and ROUGE-L scores across all difficulty levels. This is consistent with the behavior observed for BLEU and METEOR. Interestingly, the ROUGE-2 NMISS scores do not consistently outperform the standard ROUGE-2 scores across all models and levels. This deviation from the trend seen in ROUGE-1 and ROUGE-L might be attributed to the focus of ROUGE-2 on bigram overlaps, which are less sensitive to the additional contextual information.
Table 7: ROUGE-1 versus ROUGE-1 NMISS scores
ROUGE-1 versus ROUGE-1 NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.721 0.740 0.718 0.742 0.740 0.728
LOW llama2 0.201 0.712 0.282 0.342 0.717 0.421
LOW llama3 0.312 0.557 0.338 0.462 0.564 0.433
LOW mistral 0.434 0.489 0.421 0.531 0.492 0.466
LOW gemma 0.028 0.018 0.014 0.043 0.018 0.021
LOW gemma2 0.793 0.727 0.738 0.858 0.731 0.758
MID gpt4 0.552 0.529 0.530 0.603 0.531 0.539
MID llama2 0.153 0.510 0.211 0.280 0.511 0.308
MID llama3 0.356 0.605 0.384 0.558 0.610 0.488
MID mistral 0.434 0.547 0.447 0.547 0.550 0.498
MID gemma 0.017 0.018 0.015 0.049 0.022 0.026
MID gemma2 0.727 0.698 0.688 0.877 0.701 0.735
HIGH gpt4 0.585 0.603 0.582 0.647 0.604 0.611
HIGH llama2 0.156 0.477 0.211 0.301 0.485 0.330
HIGH llama3 0.271 0.432 0.295 0.491 0.452 0.401
HIGH mistral 0.364 0.537 0.396 0.499 0.539 0.455
HIGH gemma 0.084 0.084 0.076 0.111 0.091 0.095
HIGH gemma2 0.706 0.755 0.715 0.880 0.758 0.790
Table 8: ROUGE-2 versus ROUGE-2 NMISS scores
ROUGE-2 versus ROUGE-2 NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.665 0.682 0.669 0.665 0.682 0.669
LOW llama2 0.131 0.568 0.191 0.131 0.568 0.191
LOW llama3 0.218 0.449 0.243 0.223 0.450 0.245
LOW mistral 0.337 0.409 0.340 0.338 0.409 0.340
LOW gemma 0.020 0.005 0.008 0.020 0.005 0.008
LOW gemma2 0.730 0.687 0.692 0.735 0.688 0.693
MID gpt4 0.490 0.486 0.485 0.490 0.486 0.485
MID llama2 0.084 0.353 0.120 0.085 0.353 0.121
MID llama3 0.270 0.472 0.284 0.273 0.473 0.286
MID mistral 0.326 0.453 0.349 0.333 0.454 0.351
MID gemma 0.004 0.003 0.004 0.004 0.003 0.004
MID gemma2 0.662 0.637 0.632 0.665 0.638 0.633
HIGH gpt4 0.493 0.505 0.496 0.494 0.505 0.496
HIGH llama2 0.075 0.280 0.106 0.076 0.281 0.107
HIGH llama3 0.155 0.295 0.179 0.163 0.296 0.181
HIGH mistral 0.242 0.412 0.281 0.244 0.413 0.282
HIGH gemma 0.067 0.056 0.060 0.067 0.056 0.060
HIGH gemma2 0.635 0.664 0.641 0.639 0.664 0.642
Table 9: ROUGE-L versus ROUGE-L NMISS scores
ROUGE-L versus ROUGE-L NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.709 0.722 0.705 0.719 0.722 0.707
LOW llama2 0.171 0.657 0.245 0.206 0.658 0.284
LOW llama3 0.274 0.513 0.301 0.337 0.514 0.338
LOW mistral 0.390 0.454 0.385 0.440 0.456 0.406
LOW gemma 0.023 0.016 0.012 0.033 0.016 0.018
LOW gemma2 0.779 0.719 0.729 0.805 0.721 0.735
MID gpt4 0.538 0.519 0.519 0.575 0.520 0.523
MID llama2 0.129 0.460 0.180 0.165 0.461 0.214
MID llama3 0.321 0.555 0.345 0.373 0.555 0.377
MID mistral 0.377 0.500 0.398 0.413 0.502 0.413
MID gemma 0.011 0.013 0.009 0.022 0.014 0.014
MID gemma2 0.702 0.673 0.665 0.743 0.673 0.677
HIGH gpt4 0.551 0.562 0.546 0.573 0.562 0.554
HIGH llama2 0.120 0.406 0.167 0.159 0.408 0.201
HIGH llama3 0.227 0.382 0.252 0.293 0.389 0.285
HIGH mistral 0.316 0.489 0.353 0.356 0.489 0.372
HIGH gemma 0.079 0.079 0.072 0.092 0.084 0.082
HIGH gemma2 0.677 0.718 0.684 0.707 0.718 0.698
To provide a broad perspective on the various scores, Figures 5-6 illustrate the percentage of instances where NMISS outperforms standard metrics for all questions. For ROUGE-1 and ROUGE-L, NMISS consistently performs better, reflecting its ability to capture more contextually rich and detailed responses. This further highlights NMISS’s effectiveness in evaluating model performance, particularly in greater complexity scenarios.
The performance of ROUGE NMISS for precision across all question levels is as follows: for GPT-4 ROUGE NMISS outperforms ROUGE-1 by approximately 26% at the low level, 30% at the mid level, and 40% at the high level. ROUGE-L shows a similar improvement at all levels, while ROUGE-2 exhibits no notable change. For Llama2 and Llama3, ROUGE NMISS shows remarkable improvement in ROUGE-1, with gains between 80% and 85% across both low and high levels. ROUGE-L demonstrates more modest increases, ranging from 40% to 60%. For Mistral, NMISS provides more modest improvements, ranging from 4% to 65% in ROUGE-1 and similar results for ROUGE-L. Gemma2 follows a similar pattern to GPT-4, with modest improvements across all levels. At the same time, Gemma shows little to no improvement with NMISS, indicating that its issues may stem from machine-related problems rather than contextual relevance.
For recall, similar trends are observed, though with different magnitudes. This is expected, as recall is more sensitive to the length of the context, but they remain noteworthy. GPT-4 gains 18% for ROUGE-1 at mid-level questions, Llama3 improves by 11% at low and 12% at high levels, and Mistral improves by 6% at the low level. Gemma shows no change at the low level but gains 6.1% at mid and high levels.
The same patterns for the models across the question levels apply to METEOR and BLEU percentage analysis, where the difference stands in magnitude. While NMISS heavily impacts BLEU, METEOR is less pronounced.
Figure 5: Percentage of Outperforming of ROUGE NMISS scores
Figure 6: Percentage of Outperforming of BLEU AND METEOR NMISS scores
Overall, the NMISS algorithm demonstrates the greatest improvement at the highest difficulty level, where the gap between NMISS and traditional metrics like BLEU and METEOR is most pronounced. This suggests that NMISS is particularly effective in advanced scenarios that require deeper comprehension and more complex information retrieval or when the model’s predictive power is limited. Notably, with GPT-4 or Gemma2 being the top performers, they benefit the least from NMISS, indicating that answers are already highly contextual and do not need additional elaboration. In contrast, middle-range models like Llama2, Llama3, and Mistral show the most significant gains, as NMISS helps to capture their ability to provide richer contextual information. Gemma sees little to no improvement, indicating that its issues are likely machine-related rather than due to contextual relevance.
These results suggest that when models struggle to provide exact answers, they compensate by offering more detailed contextual information. Importantly, this does not result in hallucinations but rather reflects the models’ ability to maintain accuracy and completeness in the surrounding context. In such cases, even peripheral contexts can contribute to higher NMISS scores, improving the overall response quality. Conversely, where questions are more straightforward, the gap between NMISS and traditional metrics narrows. This indicates that less contextual elaboration is required, and the models perform similarly across both metrics. In summary, NMISS tends to reflect the models’ ability to provide richer, more informative answers in complex situations, which is less evident at lower difficulty levels where responses tend to be more concise and direct, reducing the gap between NMISS and traditional metrics.
6 Conclusions
In this paper, I investigate hallucinations in LLMs using a combination of detection and mitigation techniques. I evaluate a user-agent chatbot interaction with Italian health news articles through a question-answering framework based on the RAG architecture. Models including Gemma, Gemma2 GPT-4, Llama2, Llama3, and Mistral are analyzed. To enhance traditional evaluation metrics like ROUGE, METEOR, and BLEU, I introduce the NMISS scoring system, which emphasizes contextual relevance in model-generated responses. NMISS allows the identification of instances where traditional metrics undervalue contextually rich answers that deviate from the reference.
Regarding performance, Gemma2 and GPT-4 consistently outperformed other models. Gemma2 demonstrated strong results across ROUGE, METEOR, and BLEU, making it a top choice for healthcare-related chatbot applications. GPT-4 excelled in EM accuracy, highlighting its ability to generate sequences closely aligned with reference answers.
The mid-range models Llama2, Llama3, and Mistral show similar performance patterns. Although their F1 scores across all ROUGE variants are lower due to weaker precision, their recall scores are close to those of Gemma2 and GPT-4, indicating their ability to capture a significant amount of relevant content despite generating less precise outputs. Mistral maintains consistent performance across all difficulty levels, while Llama2 and Llama3 struggled with more complex queries.
NMISS application reveals that when a model struggles to provide exact answers, it often compensates by delivering richer contextual information. This does not result in hallucinations but highlights the LLM’s capacity to provide valuable information embedded in the context, even when direct answers are challenging to generate. This demonstrates the importance of context as a carrier of valuable information. NMISS consistently outperforms traditional ROUGE-1 and ROUGE-L scores across all question levels, and similar trends are observed for BLEU and METEOR.
It is important to note that high-performing models like Gemma2 and GPT-4 benefit less from NMISS, which is expected. These models already provide highly precise and contextually accurate responses. In contrast, NMISS proves particularly beneficial for mid-tier models like Llama2, Llama3, and Mistral, which often generate contextually richer responses when faced with more challenging queries. These models benefit from NMISS because it captures and rewards their ability to provide additional relevant context, even if their precision is lower. On the other hand, Gemma shows slight improvement across all metrics, suggesting that its hallucination issues are machine-intrinsic and less related to context-awareness, which NMISS is designed to evaluate.
While RAG models mitigate hallucinations by grounding outputs in external data, NMISS refines the evaluation process. It helps distinguish “false positives” — cases where traditional metrics flag hallucinations due to deviations from the reference, even though the response is contextually accurate. NMISS reduces penalties for models that provide relevant contextual information, even when their answers do not match the reference. This distinction underlines the complementary nature of RAG and NMISS: RAG acts as a tool for preventing hallucinations, while NMISS improves how I evaluate context and relevance in model output, serving as a diagnostic tool for better understanding hallucinations.
Future research could explore extending this framework to more complex datasets, tasks, and languages, further advancing hallucination mitigation and improving LLM performance in real-world applications.
Acknowledgements
This publication was produced while attending the PhD program in Economics and Business at the University of Cagliari, Cycle 38, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”.

### v3

@@
 Abstract
-I combine detection and mitigation techniques to addresses hallucinations in Large Language Models (LLMs).
-Mitigation is achieved in a question-answering Retrieval-Augmented Generation (RAG) framework while detection is obtained by introducing the Negative Missing Information Scoring System (NMISS), which accounts for contextual relevance in responses.
-While RAG mitigates hallucinations by grounding answers in external data, NMISS refines the evaluation by identifying cases where traditional metrics incorrectly flag contextually accurate responses as hallucinations.
-I use Italian health news articles as context to evaluate LLM performance.
-Results show that Gemma2 and GPT-4 outperform the other models, with GPT-4 producing answers closely aligned with reference responses.
-Mid-tier models, such as Llama2, Llama3, and Mistral benefit significantly from NMISS, highlighting their ability to provide richer contextual information.
-This combined approach offers new insights into the reduction and more accurate assessment of hallucinations in LLMs, with applications in real-world healthcare tasks and other domains.
+This work addresses hallucinations in Large Language Models (LLMs) by combining detection and mitigation strategies within a knowledge-driven evaluation framework.
+I employ a Retrieval-Augmented Generation (RAG) setup to ground model responses in structured external data, thereby improving factual reliability and helping models better follow instructions in zero-shot prompting settings.
+To enhance evaluation beyond surface-level overlap, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring system that complements traditional metrics such as BLEU, ROUGE, and METEOR by incorporating contextual relevance.
+NMISS helps distinguish between genuine hallucinations and responses that correctly infer missing information from trusted sources.
+Applied to a health-related question-answering task using Italian news articles, NMISS reveals important differences in model behavior.
+While Gemma2 and GPT-4 produce the most reference-aligned answers under standard metrics, Gemma2 also achieves the highest NMISS outperformance, suggesting that its semantically rich outputs are often under-recognized by traditional scores.
+In contrast, GPT-4 gains less from NMISS due to its strict adherence to prompt phrasing, leaving little room for contextual enhancement.
+Mid-tier models such as LLaMA2, LLaMA3, and Mistral show substantial NMISS gains when scored with BLEU and ROUGE-1, indicating their ability to provide contextually meaningful content even when traditional scores are lower.
+Overall, the RAG–NMISS framework improves hallucination detection, reveals context-aware model behavior often overlooked by traditional metrics, and supports more reliable evaluation in real-world structured knowledge tasks.
 Keywords: Large Language Models, Retrieval-Augmented Generation, Hallucinations, Healthcare, Natural Language Processing, Evaluation Metrics, Information Systems
 1 Introduction
-Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by Vaswani [52].
-Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the nuances of language.
+Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by vaswani2017attention.
+Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the features of language.
 These models perform various tasks, from basic text classification to advanced text generation, and more recent iterations can even interpret images and generate code.
 LLMs have surged in popularity due to their intuitive, prompt-based interfaces, enabling effortless interaction.
-The past few years have seen the continuous development of notable LLMs like ChatGPT [33, 1], Gemini [11], Falcon [37], Llama [49, 50, 30], and Gemma [46, 47].
-As highlighted by [15], these models excel in language comprehension [14, 16], text summarization [59], and reasoning [54, 22, 38].
+The past few years have seen the continuous development of notable LLMs like ChatGPT [OpenAI2022, achiam2023gpt], Gemini [Google2023], Falcon [penedo2023refinedweb], Llama [touvron2023llama, touvron2023llama2, meta2024llama3], and Gemma [team2024gemma, team2024gemma2].
+As highlighted by huang2023survey, these models excel in language comprehension [hendrycks2020measuring, huang2024c], text summarization [zhang2024benchmarking], and reasoning [wei2023simple, kojima2022large, qiao2022reasoning].
 As a result, their applications span various industries, such as processing electronic medical records, matching clinical trials, and drug discovery in healthcare.
-In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [35].
-However, LLMs face a significant challenge: hallucinations [12, 51].
-Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [17].
-This behavior is a serious threat since it affects the application of LLMs to real-world cases [20], as hallucinations can undermine the reliability of their outputs.
-[15] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context.
+In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [ozdemir2023quick].
+However, LLMs face a significant challenge: hallucinations [guerreiro2023hallucinations, varshneystitch].
+Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [ji2023survey].
+This behavior is a serious threat since it affects the application of LLMs to real-world cases [kaddour2023challenges], as hallucinations can undermine the reliability of their outputs.
+[huang2023survey] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context.
 Factuality hallucinations manifest as either factual inconsistencies or fabrications.
@@
 The mitigation of hallucinations has become an urgent concern due to the widespread use of LLMs in sensitive areas such as healthcare.
-Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [34], their opaque black-box nature that hides knowledge boundaries [45, 42], and their training on large datasets that may contain outdated or incorrect information [60].
-Additionally, their versatility across many tasks complicates mitigation efforts, while their strong performance can make hallucinations, especially factual fabrications, harder to detect.
-Researchers focused on various techniques to mitigate hallucinations.
-One of the most promising methods is Retrieval-Augmented Generation (RAG) [24], which incorporates external knowledge during text generation to anchor outputs in factual information.
-RAG addresses issues of accuracy and currency in LLM responses by using authoritative sources, generating verifiable answers, and enhancing user confidence across applications [21].
-Other approaches include Prompt Engineering that involves optimizing LLMs outputs with different instructions [48, 55], self-refinement through feedback and reasoning [27], where feedback on LLM outputs improves the accuracy and quality of subsequent responses, and prompt tuning, which modifies instructions during fine-tuning to improve task-specific performance [23].
-Mitigating hallucinations requires first their detection, and various metrics have emerged.
-Traditional metrics like ROUGE [26], BLEU [36], and METEOR [4] measure consistency between generated text and source text, but they often struggle with semantically relevant but superficially divergent text [5].
-To address these limitations, newer embedding-based methods such as BERTScore [58] use contextual embeddings to assess semantic similarity between sentences.
-Another approach involves ”opening the box” where token probabilities or logit outputs of LLMs are analyzed to detect hallucinations [3, 51].
-However, these techniques require access to internal model states, which is not always feasible for proprietary systems.
-This paper proposes a general framework for detecting and mitigating hallucinations, with an application to the Italian healthcare domain.
-For the mitigation part, I employ a RAG architecture to integrate external knowledge into the generation process, ensuring that the models are grounded in factual data.
-The methodology applies a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval.
-I compare a range of LLMs, including Gemma [46], Gemma 2 [47], Mistral [18], GPT-4 [1], Llama2 [50], and Llama3 [30] for the answer generation.
-This framework uses a prompting strategy to instruct models to generate answers based on the provided context.
-To assess interpretability and difficulty, I categorize 100 questions into three levels of understanding: low, mid, and high.
-For the detection strategy, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring framework designed to be applied alongside existing metrics to integrate contextual relevance.
-By analyzing residual words from machine-generated answers after comparison with human responses, NMISS refines traditional metrics by highlighting cases where they may unfairly penalize performance despite the enhanced contextual relevance of the responses.
-Results show that Gemma2 and GPT-4 achieve the highest overall performance in Italian healthcare-related question-answering tasks across all difficulty levels, while Gemma exhibits the highest rate of hallucinations.
-The NMISS scoring system provides a more accurate assessment of LLM performance, especially in high-difficulty tasks for mid-tier models like Llama2, Llama3, and Mistral, which show notable improvements compared to traditional metrics.
-Additionally, NMISS recognizes the value of contextual information provided by models in cases where they fail to deliver exact matches to expected answers in difficult tasks.
-By focusing on the relevance and coherence of the provided context, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they do not fully align with the reference answers.
-This paper is organized as follows: Section 2 reviews the relevant literature on hallucination detection and mitigation.
-Section 3 details the methodology, highlighting the contributions of NMISS.
-Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 presents my findings.
-Finally, Section 6 summarizes contributions and outlines future research directions.
+Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [ouyang2022training], their opaque black-box nature that hides knowledge boundaries [sun2023moss, ren2023investigating], and their training on large datasets that may contain outdated or incorrect information [zhang2023siren].
+Additionally, their versatility across many tasks complicates mitigation efforts, while their performance makes hallucinations, especially factual fabrications, harder to detect.
+Researchers have explored several strategies to mitigate hallucinations.
+One promising approach is Retrieval-Augmented Generation (RAG) [lewis2020rag], which incorporates external knowledge during generation to ground outputs in factual information.
+RAG improves both the accuracy and timeliness of model responses by leveraging authoritative sources, producing verifiable answers, and increasing user confidence [kang2023ever].
+Other techniques include prompt engineering, which optimizes outputs via instruction design [tonmoy2024comprehensive, white2023prompt]; self-refinement, where feedback on prior outputs enhances subsequent responses [madaan2023selfrefineiterativerefinementselffeedback]; and prompt tuning, which adapts instruction embeddings during fine-tuning to improve task-specific performance [lester-etal-2021-power].
+Detecting hallucinations is a prerequisite for effective mitigation.
+Traditional metrics such as ROUGE [lin2004rouge], BLEU [papineni2002bleu], and METEOR [banerjee2005meteor] measure lexical overlap between generated and reference text but often fail to capture semantic equivalence [celikyilmaz2020evaluation].
+More recent embedding-based metrics, such as BERTScore [zhang2019bertscore], use contextual embeddings to assess semantic similarity.
+Another direction involves “opening the box”—analyzing token probabilities or logit outputs to detect hallucinations [azaria2023internal, varshneystitch]—though this approach requires access to internal model states, which is not feasible for proprietary systems.
+This work proposes a general, domain-independent framework for hallucination detection and mitigation in LLMs.
+For mitigation, a RAG architecture integrates external knowledge into the generation process, ensuring that outputs remain grounded in factual information.
+The methodology employs a multilingual Sentence Transformer model [reimers-2019-sentence-bert, reimers-2020-multilingual-sentence-bert] to construct and query a vector database for contextual retrieval.
+A range of LLMs, specifically Gemma [team2024gemma], Gemma-2 [team2024gemma2], Mistral [jiang2023mistral], GPT-4 [achiam2023gpt], Llama-2 [touvron2023llama2], and Llama-3 [meta2024llama3] are compared using a zero-shot prompting strategy that instructs models to generate answers based on retrieved context.
+To evaluate interpretability and task complexity, 100 questions are categorized into three difficulty levels: low, medium, and high.
+For detection, this work introduces the Negative Missing Information Scoring System (NMISS), a novel evaluation layer that extends traditional metrics by incorporating contextual relevance.
+Unlike prior metrics, which assess only lexical or semantic overlap with reference answers, NMISS leverages context from external sources to distinguish genuine hallucinations from contextually valid elaborations.
+Specifically, it examines lexical items present in model outputs but absent from human references, evaluating whether these additions are substantiated by the provided context.
+Through this mechanism, NMISS refines traditional scoring methods to more accurately reflect whether supplementary content contributes relevant information or constitutes hallucination.
+Under the RAG-based evaluation, Gemma-2 and GPT-4 achieve the highest overall performance across all difficulty levels, demonstrating strong grounding and retrieval effectiveness, while Gemma exhibits the highest hallucination rate.
+NMISS, in turn, provides a more reliable assessment of model behavior, particularly in high-difficulty cases for mid-tier models such as Llama-2, Llama-3, and Mistral, which show notable improvements under this context-aware metric.
+Interestingly, models with more verbose generation styles tend to benefit from NMISS, as their outputs often retrieve and reuse lexical elements directly from the provided context.
+This might reflect computational constraints or limited abstraction capacity.
+Nevertheless, Gemma-2 also benefits from this behavior, suggesting that contextual retrieval contributes positively to factual grounding rather than mere verbosity.
+By emphasizing contextual relevance and coherence over strict lexical matching, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they diverge from the reference answers.
+The main contributions of this paper are as follows: (i) it proposes NMISS, a novel scoring framework for hallucination detection that extends standard evaluation metrics by incorporating contextual relevance, thereby enabling a more accurate distinction between hallucinations and contextually valid elaborations.
+The framework is general-purpose and applicable beyond the healthcare domain; (ii) it develops a RAG–based evaluation framework that grounds LLM outputs in structured external knowledge, demonstrating that retrieval mitigates hallucinations and enhances factual grounding; (iii) it presents a comprehensive evaluation of multiple LLMs across varying levels of reasoning difficulty, showing that NMISS exposes performance differences overlooked by traditional metrics; and (iv) it introduces a new dataset comprising 100 Italian healthcare questions annotated by reasoning complexity, supporting domain-specific benchmarking and multilingual question-answering evaluation.
+The work is organized into five sections.
+Section 2 reviews the relevant literature on hallucination detection and mitigation.
+Section 3 presents the proposed methodology and highlights the contributions of NMISS.
+Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 reports the findings.
+Finally, Section 6 concludes the study by summarizing the main contributions and outlining future research directions.
 2 Related Works
+This section reviews recent research on hallucination management in LLMs, organized along two main dimensions.
+First, it distinguishes between detection techniques, which aim to identify hallucinated content, and mitigation techniques, which seek to reduce hallucinations during text generation.
+Second, within mitigation, existing approaches are further classified by their resource requirements: resource-intensive strategies that rely on model internals or external data, and zero-resource strategies that operate in black-box settings and often combine elements of both detection and mitigation.
+This organization clarifies the conceptual landscape of hallucination research and situates the contributions of this work within it.
 2.1 Hallucination Detection Techniques
-Hallucinations in LLMs have been widely studied.
-[15] provides a detailed overview of strategies to detect and mitigate hallucinations, offering insight into their occurrence in current LLMs architectures.
-Ensuring LLM’s trustworthiness depends on early detection, as emphasized by [51], who highlights the importance of proactive measures during text generation.
-[60] categorize detection methods into three main types.
-First, logit-based estimation, which uses token-level probabilities or entropy from the model’s logits [39].
-Second, verbalized confidence estimation, where LLMs are asked to self-rate the confidence in their responses.
-Third, consistency-based estimation involves generating multiple candidate outputs and comparing them to detect logical inconsistencies, mimicking human reasoning to refine answers [53, 44, 61].
-Furthermore, [19] categorize hallucination detection into white-box, grey-box, and black-box techniques.
-White-box methods analyze the model’s internal workings to identify when incorrect statements are generated.
-Grey-box approaches examine logit outputs to find areas of low confidence in the output.
-Black-box techniques, which only require the model’s final answers, are simpler to apply since they don’t require access to internal states.
-Recent studies introduce SAPLMA [3], which detects hallucinations by using a classifier trained to predict the truthfulness of statements based on LLMs’ hidden layer activations.
-This approach outperforms traditional methods that rely solely on sentence probabilities.
-[51] propose an iterative process that detects potential hallucinations by identifying tokens with low normalized probabilities.
-The system then retrieves relevant information to correct these errors, effectively reducing hallucinations in GPT-3.5.
+Hallucinations in LLMs have been widely studied in recent years.
+huang2023survey provide a comprehensive overview of strategies for detecting and mitigating hallucinations, offering valuable insight into how they emerge in current LLM architectures.
+Ensuring the trustworthiness of LLMs depends critically on early detection, as emphasized by varshneystitch, who highlight the importance of proactive identification during text generation.
+zhang2023siren categorize detection methods into three main types.
+First, logit-based estimation, which relies on token-level probabilities or entropy derived from model logits [quevedo2024detecting].
+Second, verbalized confidence estimation, in which LLMs are prompted to self-rate their confidence in generated responses.
+Third, consistency-based estimation, which compares multiple generated outputs to identify logical inconsistencies, thereby mimicking human reasoning to refine responses [wang2022self, shi2023replug, zhao2023verify].
+Complementary to this taxonomy, jiang2024hallucination classify hallucination detection methods into white-box, grey-box, and black-box approaches.
+White-box methods analyze internal model activations to pinpoint when incorrect statements are produced, while grey-box methods examine logit-level confidence to detect uncertainty.
+In contrast, black-box methods rely solely on model outputs and are easier to apply, since they do not require access to internal states.
+Recent studies have also introduced specialized architectures.
+SAPLMA [azaria2023internal], for example, detects hallucinations using a classifier trained on hidden layer activations to predict the truthfulness of statements, outperforming traditional probability-based baselines.
+Likewise, varshneystitch propose an iterative detection process that identifies low-probability tokens and retrieves external evidence to correct potential errors, effectively reducing hallucinations.
+Alongside model-internal techniques, several surface-level metrics have been employed to detect hallucinations based on output similarity.
+Traditional reference-based metrics such as ROUGE [lin2004rouge], BLEU [papineni2002bleu], and METEOR [banerjee2005meteor] measure lexical overlap between generated and reference text, but often fail to capture semantically valid variations or long-range dependencies [celikyilmaz2020evaluation].
+BERTScore [zhang2019bertscore], which leverages contextual embeddings from BERT [devlin2018bert], improves upon these by evaluating token-level semantic similarity.
+However, even BERTScore does not account for whether additional, unmatched content is contextually justified.
+This study differs from prior work by introducing the Negative Missing Information Scoring System (NMISS), a context-aware scoring framework that enhances existing metrics through contextual alignment, without requiring access to model internals or probability distributions.
 2.2 Resource-Intensive Mitigation Strategies
-In addition to detection, various strategies have been developed to mitigate hallucinations in LLMs.
-One common approach is Knowledge Augmentation, such as Knowledge Injection (KI).
-[29] enhance prompts with information from knowledge graphs, such as adding details like location names for review tasks.
-This significantly reduces hallucinations and improves factual accuracy, with KI-LLMs producing more correct assertions and outperforming larger models in response quality.
-[31] take a direct approach by training T5 models on knowledge graph triples from Wikidata.
-By masking key entities, the model learns and generalizes knowledge effectively.
-SKILL, after fine-tuning, outperforms standard T5 models on question-answering tasks, proving the effectiveness of structured data infusion.
-Another key mitigation strategy is retrieval-based frameworks.
-[25] combines a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework.
-This integration enhances the accuracy and reliability of generated content.
-Similarly, [13] apply retrieval techniques in biomedical contexts, incorporating sources like UMLS and Wikipedia to generate more factual summaries, which are crucial for healthcare applications.
-[62] introduce a self-supervision framework that assigns risk scores, called POLAR scores, to LLM responses as indicators of hallucination risk.
-Their approach uses Pareto optimization to balance supervision from multiple sources.
-The process involves gathering LLM responses, training a probabilistic function $h$ to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support.
-Experimental results show significantly reduced hallucination rates across tasks like question answering and text generation.
-[57] propose the MixAlign framework, which aligns user queries with external knowledge to resolve ambiguities.
-This improves response accuracy and reduces hallucinations, with results showing a sharp increase in correct answers and a decrease in hallucinations.
+Beyond detection, a range of strategies have been developed to mitigate hallucinations in LLMs, often requiring computational or data resources.
+One prominent category is Knowledge Augmentation, including techniques such as Knowledge Injection (KI).
+martino2023knowledge enhance prompts with information derived from knowledge graphs, such as adding geographical details for review tasks.
+This approach significantly reduces hallucinations and improves factual accuracy, with KI-enhanced LLMs producing more correct assertions and outperforming larger baseline models in response quality.
+moiseev2022skill adopt a more direct approach by training T5 models on Wikidata triple knowledge graphs.
+By masking key entities during training, the model learns to infer and generalize structured knowledge effectively.
+Their model outperforms standard T5 variants on question-answering benchmarks, demonstrating the benefits of structured data infusion.
+Another major line of research involves retrieval-based frameworks.
+lewis2021improving combine a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework, thereby improving the accuracy and factual grounding of generated content.
+Similarly, guo2024retrieval apply retrieval augmentation in biomedical contexts, using external sources to produce more accurate and verifiable summaries, which are crucial for healthcare applications.
+A different approach, introduced by zhao2023automatic, is a self-supervision framework that assigns risk scores, termed POLAR scores, to LLM responses as indicators of hallucination likelihood.
+Their method employs Pareto optimization to balance supervision signals from multiple sources.
+The process involves collecting model responses, training a probabilistic function to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support.
+Experimental results demonstrate substantial reductions in hallucination rates across various tasks, including question answering and text generation.
+In a related effort, zhang2024knowledgealignmentproblembridging propose the MixAlign framework, which aligns user queries with external knowledge sources to resolve semantic ambiguities.
+This method improves both response accuracy and factual consistency, yielding a marked increase in correct answers and a reduction in hallucination frequency.
+Unlike prior resource-intensive frameworks, this study introduces a lightweight mitigation approach that integrates zero-shot prompt engineering with RAG for factual grounding in the healthcare domain.
 2.3 Zero-Resource Hallucination Strategies
 The strategies discussed in Section 2.2 are resource-intensive.
-In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited resources, relying only on model outputs without needing internal access or external data.
-Many of these techniques not only detect hallucinations but also help mitigate them by refining responses based on internal consistency assessments.
-One approach is through self-consistency techniques, which compare multiple responses generated by the LLM itself.
-[28] propose a zero-resource method that generates multiple responses from GPT-3 and compares them to a reference answer to assess factual consistency.
-This method detects and mitigate hallucinations using low-level metrics and outperforms traditional grey-box and black-box methods.
-Another approach is self-rating techniques, as proposed by [6].
-The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers.
-RelD is trained by converting a regression task into a binary classification task, aligning with human intuition.
+In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited computational capacity, relying solely on model outputs without requiring internal access or external training data.
+Many of these techniques not only detect hallucinations but also mitigate them by refining responses through internal consistency checks.
+One common approach involves self-consistency techniques, which compare multiple responses generated by the LLM itself.
+manakul2023selfcheckgpt propose a zero-resource method that samples multiple outputs from GPT-3 and compares them to a reference answer to assess factual consistency.
+This method detects and mitigates hallucinations using low-level statistical metrics and outperforms traditional grey-box and black-box baselines.
+Another class of methods includes self-rating techniques, as proposed by chen2023hallucination.
+The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers by converting a regression task into a binary classification problem, aligning with human judgment.
 By flagging low-confidence responses, it serves as an indirect mitigation mechanism.
-Experiments across various datasets and models (e.g., BLOOM [56]) show RelD’s effectiveness, validated by both automatic metrics and human-in-the-loop evaluations.
-Similarly, [2] addresses hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources.
-DQs involve asking the model a binary question multiple times to calculate a probability.
-In contrast, IQs ask about the authors or content of the reference and check for consistency across responses.
-Although primarily designed for detection, these methods can also reduce hallucinations by discarding low-confidence responses.
-Results show that IQs generally perform better than DQs in detecting hallucinations, though performance varies among models like GPT-4, GPT-3, and Llama 2.
-Finally, [32] explore self-contradiction techniques as a factor in hallucinations.
-They introduce a prompting framework to detect and mitigate contradictions within the same response, relying solely on the model’s logical reasoning capabilities.
-This framework includes functions for gathering contextual information, generating sentences, identifying contradictions, and refining sentences.
-Across various LLMs, this approach achieves an 80% F1 score.
-2.4 Metrics for Evaluating Hallucinations
-Traditional metrics to detect or manage hallucinations [5] are notably ROUGE [26], BLEU [36], and METEOR [4].
-These indicators measure the consistency between the generated text and the source text.
-However, a drawback of these simple metrics is that they do not accurately capture distance word-dependencies or if the two answers diverge in lexicon but have equal meaning.
-BERTScore overcomes this issue [58] by leveraging pre-trained contextual embeddings from BERT [9] and matches words in candidate and reference sentences by cosine similarity, correlating well with human judgments.
-2.5 Contributions
-Despite significant progress in hallucination detection and mitigation, key challenges persist.
-Many existing methods, such as knowledge injection or retrieval, are resource-intensive, limiting their practicality in certain applications.
-Additionally, traditional evaluation metrics often fail to account for contextual accuracy, unfairly penalizing semantically correct responses that differ lexically from reference answers.
-This paper addresses these gaps in several steps. The key contributions of this work are as follows:
-(1) A combined approach for hallucination detection and mitigation in LLMs.
-I develop a RAG-based architecture to integrate external knowledge into the generation process, ensuring that models are grounded in factual data.
-The methodology leverages a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval.
-This framework includes a prompting strategy to instruct models to generate answers based on the provided context, enabling a comparative evaluation of various LLMs using this approach.
-(2) Comprehensive analysis of hallucinations across reasoning levels.
-To evaluate hallucination tendencies in LLMs, I analyze their performance across progressively challenging levels of reasoning, categorized in low, mid and high levels.
-This approach enables an assessment of how hallucination rates and mitigation effectiveness vary as tasks demand increasing degrees of logical complexity and understanding.
-The progressive evaluation highlights how model reliability and factual grounding are affected under escalating reasoning challenges, providing insights into their strengths and vulnerabilities.
-(3) Development of a detection scoring system.
-I introduce the Negative Missing Information Scoring System (NMISS), a novel framework that fine-grains traditional metrics by incorporating contextual relevance.
-By analyzing residual words from machine-generated answers after comparison with human responses, NMISS highlights cases where traditional metrics may penalize models despite the enhanced contextual relevance of their answers.
+Experiments across several datasets and models (e.g., BLOOM [workshop2022bloom]) demonstrate RelD’s effectiveness, validated through both automated metrics and human evaluations.
+Similarly, agrawal2023language address hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources.
+DQs repeatedly ask binary questions to estimate response probability, whereas IQs query contextual details such as authors or publication content to test internal consistency.
+Although primarily designed for detection, these approaches can also mitigate hallucinations by filtering out low-confidence responses.
+Results show that IQs generally outperform DQs, though effectiveness varies among models such as GPT-4, GPT-3, and Llama-2.
+Finally, mundler2023self explore self-contradiction techniques, which identify hallucinations by detecting logical inconsistencies within a single response.
+Their prompting framework relies entirely on the model’s internal reasoning capabilities and includes procedures for gathering contextual information, generating candidate sentences, identifying contradictions, and refining responses.
+Across several LLMs, this method achieves an F1 score of approximately 80%.
+In contrast to self-referential, zero-resource techniques, this work introduces a context-aware hallucination detection mechanism, NMISS, which evaluates generated outputs without relying on self-consistency signals or model confidence estimation.
+The proposed hybrid approach remains lightweight, requiring neither fine-tuning nor model access, while enhancing factual grounding through the integration of retrieved evidence.
 3 Methodology
 3.1 RAG Architecture
-RAG models combine the strengths of retrieval-based and generative language models to address issues like hallucinations and factual inaccuracies by integrating a retrieval mechanism to ground responses in relevant information [24].
+RAG models combine the strengths of retrieval-based and generative language models to address issues such as hallucinations and factual inaccuracies by integrating a retrieval mechanism that grounds responses in relevant information [lewis2020rag].
 I formally define the problem as follows.
 Let $q$ represent the query, $c$ the context information retrieved from external sources, and $r$ the output sequence generated by the model.
-The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-K set of documents $k$ based on the query $q$, and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
+The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-$K$ set of documents $k$ based on the query $q$; and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, the retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
 $p_{\text{RAG}}(r\mid q)=\sum_{c\in k(p(\cdot\mid q))}p_{\eta}(c\mid q)\prod_{i=1}^{N}p_{\theta}\left(r_{i}\mid q,c,r_{1:i-1}\right)$ (1)
-This RAG-sequence approach ensures that the response remains consistent by using the same retrieved document to generate each token, maintaining coherence throughout the sequence.
-Here, $N$ represents the length of the output sequence.
-This architecture is particularly advantageous for domains like healthcare, where factual accuracy is critical.
-By leveraging retrieval from medical documents, guidelines, or research articles, the RAG model ensures that responses are accurate, minimizing the risks associated with misinformation [10].
-In this paper, I employ the retrieval and generation models separately and combine them for inference; the workflow is detailed in Section 4.2.
+This implementation follows the RAG-Sequence variant proposed by lewis2020rag, in which a single retrieved context is used to generate the entire response.
+Unlike the RAG-Token approach, which may retrieve a different document for each generated token, the former maintains global coherence and is computationally more efficient.
+Given the goal of ensuring factual grounding and semantic consistency in healthcare-related question answering, this variant is particularly suitable for the present work.
+In this study, the retrieval and generation components are implemented separately and combined during inference; the complete workflow is described in Section 4.3.
 3.2 Evaluation Metrics
-The metrics described in this section are used to evaluate the quality of responses generated by the system, referred to as the System Summary (SS), against the ground truth responses, referred to as the Reference Summary (RS), which are human-generated responses in this study.
-Both SS and RS are treated as sets of n-grams for all metrics.
-Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$, where $s_{i}$ is an n-gram in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$, where $r_{j}$ is an n-gram in the RS.
-ROUGE [26] evaluates responses based on precision, recall, and F1-score.
-The text is pre-processed before calculating these scores [43].
-More formally, ROUGE-N measures the overlap of n-grams between the SS and RS.
-The precision, recall, and F1-score for ROUGE-N are defined as follows:
+This section describes the metrics employed to evaluate the quality of the responses generated by the system, referred to as the System Summary (SS), against the ground-truth responses, referred to as the Reference Summary (RS), which in this study correspond to human-authored answers.
+The evaluation adopts multiple complementary metrics, including ROUGE, BLEU, METEOR, BERTScore, and NMISS, which jointly assess lexical, semantic, and contextual alignment between SS and RS.
+For all metrics, both SS and RS are represented as sets of $n$-grams.
+Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$ denote the set of $n$-grams in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$ denote the set of $n$-grams in the RS.
+ROUGE [lin2004rouge] evaluates textual overlap in terms of precision, recall, and F1-score, after standard text preprocessing [sarkar2019text].
+More formally, ROUGE-$N$ measures the $n$-gram overlap between the SS and RS.
+The corresponding precision, recall, and F1-score are defined as:
 $\text{Precision}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{SS}}|}$ (2)
 $\text{Recall}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{RS}}|}$ (3)
-[…]
-where $S_{\text{SS}}$ is the set of n-grams in the SS, $S_{\text{RS}}$ is the set of n-grams in the RS, $|S_{\text{SS}}\cap S_{\text{RS}}|$ is the cardinality of the intersection of $S_{\text{SS}}$ and $S_{\text{RS}}$, $|S_{\text{SS}}|$ is the cardinality of the set $S_{\text{SS}}$, and $|S_{\text{RS}}|$ is the cardinality of the set $S_{\text{RS}}$.
-Precision measures the accuracy of the generated text, while recall evaluates the coverage of the information.
-Low precision suggests hallucination, while low recall indicates a coverage issue.
-Another variant used here is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS.
-The LCS represents the longest subsequence in both SS and RS while preserving word order, though the words do not need to appear consecutively [7].
-BLEU [36] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers.
+$\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
+where $|S_{\text{SS}}\cap S_{\text{RS}}|$ denotes the number of overlapping $n$-grams, $|S_{\text{SS}}|$ the total number of $n$-grams in the SS, and $|S_{\text{RS}}|$ the total number of $n$-grams in the RS.
+Precision measures the factual accuracy of the generated text, while recall quantifies the coverage of the reference information.
+Accordingly, low precision suggests the presence of hallucinations, whereas low recall indicates the presence of missing or incomplete content.
+A widely used variant is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS.
+The LCS represents the longest ordered sequence of words appearing in both summaries, without requiring consecutive word matching [cormen2022introduction].
+ROUGE-L thus captures sentence-level fluency and structural similarity, complementing the $n$-gram–based metrics.
+BLEU [papineni2002bleu] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers.
 BLEU accounts for variability in word choice and word order found in different, yet valid, translations of the same source sentence.
-It quantifies the overlap by measuring how many n-grams from the SS are present in the RS.
+It quantifies the overlap by measuring the number of $n$-grams from the SS that are present in the RS.
 The core of the BLEU metric is the modified $n$-gram precision, which addresses the over-generation problem by treating each reference word as exhausted once a matching candidate word is found.
@@
 where $b$ is the length of the candidate translation (SS), and $r$ is the effective reference length (RS).
-The final BLEU score is computed by taking the geometric mean of the modified n-gram precisions (up to length $N$, typically set to 4), and multiplying by the brevity penalty:
+The final BLEU score is computed by taking the geometric mean of the modified $n$-gram precisions, up to length $N$ and typically set to 4, and multiplying by the brevity penalty:
 $\text{BLEU}=\text{BP}\cdot\exp\left(\frac{1}{N}\sum_{n=1}^{N}\log p_{n}\right)$ (7)
-METEOR [4] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms.
-METEOR computes the score by comparing the SS to each RS and reporting the highest score.
-Additionally, METEOR introduces a Discontiguous Matches Penalty (DP), based on the number of matched chunks.
-Let $T$ represent the number of contiguous chunks of matched unigrams between SS and RS.
-For a set of matched unigrams $V=\{v_{1},v_{2},\dots,v_{k}\}$, a chunk is defined as a maximal sequence of consecutive matches $v_{i},v_{i+1},\dots,v_{i+t}$, such that $v_{i+t}=v_{i}+t$ for all $t\geq 0$.
-The function $T(V,S_{\text{SS}})$, which counts the number of contiguous chunks, is then formalized as:
-$T(V,S_{\text{SS}})=\left|\{\{v_{i},v_{i+1},\dots,v_{i+t}\}\subseteq V\mid v_{i+t}=v_{i}+t\}\right|$ (8)
-This function returns the number of contiguous chunks of matched unigrams in $S_{\text{SS}}$. The Discontiguous Matches Penalty (DP) is then computed as:
-$\text{DP}=0.5\cdot\left(\frac{T(V,S_{\text{SS}})}{|V|}\right)$ (9)
-where $|V|$ is the number of matched unigrams.
-The penalty increases as the number of chunks grows, penalizing fragmented matches.
-The final METEOR score is adjusted by applying DP to the $F_{\text{mean}}$ score:
-$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{DP})$ (10)
-BERTSCORE is an evaluation metric for text generation that leverages pre-trained BERT embeddings [58].
-Unlike traditional metrics that rely on exact n-gram matches, BERTSCORE measures the semantic similarity between tokens in the System Summary (SS) and Reference Summary (RS) by comparing their contextual embeddings.
+METEOR [banerjee2005meteor] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms.
+For each SS–RS pair, METEOR finds an alignment of unigrams that maximizes the number of matches subject to one-to-one mapping and minimal crossing.
+Precision ($P$) and recall ($R$) are computed as the fractions of matched unigrams in the SS and RS, respectively, and combined through a weighted harmonic mean:
+$F_{\text{mean}}=\frac{P\cdot R}{\alpha P+(1-\alpha)R},$
+where $\alpha$ is typically set to 0.9.
+To capture word order and fragmentation, METEOR groups matched unigrams into the fewest possible number of contiguous chunks such that the words in each chunk are adjacent and in the same order in both SS and RS.
+Let $T$ denote the number of such chunks, and $|V|$ the total number of matched unigrams.
+The fragmentation penalty is then computed as:
+$\text{Penalty}=0.5\left(\frac{T}{|V|}\right)^{3}.$
+The final METEOR score is given by:
+$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Penalty}).$
+This formulation penalizes disordered or fragmented alignments, ensuring that longer, well-ordered matches contribute more to the final score.
+BERTScore is an evaluation metric for text generation that leverages pre-trained BERT embeddings [zhang2019bertscore].
+Unlike traditional metrics that rely on exact $n$-gram matches, BERTScore measures the semantic similarity between tokens in the SS and the RS by comparing their contextual embeddings.
 This method captures the semantic meaning even when the exact words differ.
-BERTSCORE generates contextual embeddings for the tokens in both the SS and RS, where each embedding reflects the surrounding context of the word.
+It generates contextual embeddings for tokens in both the SS and RS, where each embedding reflects the surrounding context of the word.
 The similarity between tokens is computed using cosine similarity.
 For each token in the RS, BERTScore identifies the most similar token in the SS, and vice versa, to compute precision and recall.
-Given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token.
+Formally, given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token.
 The similarity between two embeddings is computed using cosine similarity.
@@
 $EM=\begin{cases}1&\text{if }SS=RS\\ 0&\text{if }SS\neq RS\end{cases}$ (11)
-3.3 NMISS
+3.3 Negative Missing Information System Scoring
 3.3.1 Rationale
-The metrics discussed above evaluate the quality of text summaries and machine-generated responses by comparing them to reference answers.
-While effective, this method overlooks the possibility that non-matching words or n-grams might still be valuable if they align with the question’s broader context.
-This observation prompts an investigation into whether the words or phrases that do not match the reference could still be relevant compared to the broader context in which the response is generated.
-LLMs often produce responses beyond the strict instructions, frequently adding extra content, as shown in Table 1.
-This additional content can manifest in two ways: as hallucinations, where the model generates information that is factually inaccurate or irrelevant, or as meaningful elaborations, where the model extends the response in a contextually relevant manner.
+The metrics discussed above evaluate the quality of machine-generated responses by measuring their lexical or semantic overlap with reference answers.
+While effective, these approaches overlook the fact that non-matching words or $n$-grams may still carry meaningful information if they align with the broader context of the question or source material.
+In practice, LLMs often extend their responses beyond the explicit instructions, adding supplementary content, as illustrated in Table 1.
+Such additions can take two forms: i) hallucinations, where the model introduces factually incorrect or irrelevant information, and ii) contextually relevant elaborations, where the model enriches the answer by incorporating information consistent with the provided context, as in RAG-based generation.
+Recognizing this distinction is essential for a more accurate evaluation of model behavior.
+By explicitly accounting for non-matching elements, it becomes possible to distinguish between genuinely hallucinated content and valuable, contextually appropriate extensions.
+The NMISS captures the potential informational value of these unmatched words by assessing their alignment with the contextual evidence provided.
+This rationale acknowledges that additional content in LLM outputs is not inherently erroneous but can instead reflect a form of reasoning or contextual expansion that traditional metrics fail to reward.
 Table 1: Example of Question, Context, Reference, and System Answer
@@
 However, based on the context, it is noted that the Relaunch Decree emphasizes the importance of strict behaviors to minimize impacts at all levels, such as social distancing, mask use, and hand hygiene, which are generally useful measures for preventing disease transmission in general, not just the discussed epidemic.
-By accounting for these non-matching elements, I can more effectively evaluate the performance of LLMs, distinguishing between irrelevant or hallucinated information and valuable, contextually appropriate elaborations.
-This approach offers a more nuanced understanding of the model’s output.
-The NMISS scoring system was developed to capture the potential value in these unmatched words by evaluating whether they align with the contextual information provided.
-This rationale stems from the observation that LLMs often do not strictly adhere to the given instructions, incorporating additional content into their answers.
-This content may either be a hallucination or a contextually relevant elaboration that enriches the response.
 3.3.2 Formulation
-Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer.
+Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer.
 Define the set of matching tokens between system-generated answers and reference answers as $\mathcal{P}(s_{i},r_{i})$, and let $\mathcal{N}(s_{i},r_{i})=s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ represent the tokens in $s_{i}$ that do not match $r_{i}$.
 Finally, let $\mathcal{A}(s_{i},c_{i})=\mathcal{N}(s_{i},r_{i})\cap c_{i}$, which represents the set of unmatched tokens in $s_{i}$ that match the context $c_{i}$.
-I aim to define a function
+The function is defined as:
 $f_{NMISS}:\mathcal{S}\times\mathcal{R}\times\mathcal{C}\rightarrow\mathbb{R}$ (12)
-that evaluates how well system-generated answers align with both reference answers and the context.
-First, I define a reference-based evaluation function $f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$, and a context-based evaluation function $f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$, which computes the score based on tokens matching the context.
-The NMISS evaluation function is defined as:
+which evaluates how well system-generated answers align with both reference answers and contextual information.
+The reference-based evaluation function is defined as follows:
+$f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$ (13)
+while the context-based evaluation function, which computes the score based on tokens matching the context, is defined as:
+$f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$ (14)
+The NMISS evaluation function is given by:
 $f_{\text{NMISS},i}(s_{i},r_{i},c_{i})=\max\left(f_{\text{ref},i}(s_{i},r_{i}),\frac{\lambda_{1}\cdot f_{\text{ref},i}(s_{i},r_{i})+\lambda_{2}\cdot f_{\text{cxt},i}(s_{i},c_{i})}{\lambda_{1}+\lambda_{2}}\right)$ (15)
@@
 $\lambda_{1}$ and $\lambda_{2}$ represent the number of matching reference and context matches tokens, respectively.
-To safeguard the integrity of standard scoring metrics, NMISS implements two key mechanisms: a weighted scheme and a maximum function.
-These mechanisms are essential for balancing reference-based and context-based contributions, addressing challenges such as score underestimation or overestimation when processing contextual information.
-The weighted scheme ensures a fair combination of reference-based and context-based scoring, preventing a disproportionate reliance on one over the other.
-Specifically, $\lambda_{1}$ and $\lambda_{2}$ control the relative influence of reference overlap and contextual alignment.
-This mechanism helps NMISS capture instances where a system-generated response extends beyond the reference but remains faithful to the context, distinguishing it from hallucinated content.
-The maximum function further enhances robustness by preventing NMISS from being unfairly penalized when context length inflates the denominator.
-In recall-sensitive metrics, an inflated denominator—such as the total number of n-grams in the context—could lower the score even if the additional content is meaningful.
-The formulation in Equation 16 ensures that context-based contributions only positively influence the score when they genuinely add value.
-In other words, if context matching introduces more noise than useful content, NMISS defaults to the reference-based score.
-This design prevents cases where large but irrelevant contexts artificially suppress the metric.
-This approach is particularly advantageous in tasks like question-answering, where the model’s response might not strictly match the reference but still be valid given the context.
-By integrating these mechanisms, NMISS ensures coherence in scoring, making it a more flexible and context-aware evaluation metric.
-The detailed implementation is outlined in Algorithm 1.
+To preserve the integrity of standard scoring metrics, NMISS introduces two key mechanisms: a weighted scheme and a maximum function.
+These mechanisms jointly balance reference-based and context-based contributions, mitigating potential distortions such as score inflation or deflation when contextual information is incorporated.
+The weighted scheme harmonizes the relative influence of reference overlap and contextual alignment.
+Because contextual passages are typically much longer and information-rich than reference responses, a simple averaging would overemphasize the contextual component.
+To address this, the weights $\lambda_{1}$ and $\lambda_{2}$ regulate their respective contributions, ensuring that neither dominates the final score.
+This mechanism enables NMISS to recognize cases where a system-generated response extends beyond the reference while remaining coherent and faithful to the retrieved evidence.
+The maximum function serves as a safeguard against disproportionate score drops caused by denominator inflation in recall-sensitive formulations.
+This situation often arises with verbose models that rely on contextual paraphrasing when uncertain.
+Although such verbosity may not introduce hallucinations, it can generate noise—information that adds limited value relative to the reference.
+By applying a floor, NMISS reverts to the reference-based score whenever the context-based contribution disproportionately lowers the metric due to a large $\lambda_{2}$ term in the denominator.
+This ensures rewarding genuine contextual relevance while preventing technical artifacts from suppressing the overall score.
+This scoring framework is particularly advantageous in question-answering tasks, where model responses may diverge lexically from the reference yet remain contextually valid.
+The detailed implementation is provided in Algorithm 1.
 Algorithm 1 NMISS Algorithm
-1:=$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, the set of reference answers
-2:$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, the set of system-generated answers
-3:$\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$, the set of context texts
-4:=$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, the set of NMISS scores for a metric $\mathcal{X}$
+[…]
+[…]
+[…]
+[…]
 5:
@@
 7:  Stage 1: Reference-Based Metric Calculation
-8:     $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$triangleright Matched tokens
-9:     $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$triangleright Unmatched tokens
+8:  $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$ $\triangleright$ Matched tokens
+9:  $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ $\triangleright$ Unmatched tokens
 10:  $f_{\text{ref},i}(s_{i},r_{i})\leftarrow\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$
@@
 12:  Stage 2: Context-Based Metric Calculation
-13:     $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$triangleright Context-matched tokens
+13:  $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$ $\triangleright$ Context-matched tokens
 14:  $f_{\text{cxt},i}(s_{i},c_{i})\leftarrow\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$
@@
 20:end for
+To formally evaluate NMISS outperformance, let $\mathcal{H}=\{h_{1},h_{2},\dots,h_{n}\}$ denote the set of hallucination flags associated with the model-generated responses $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$.
+Each flag $h_{i}\in\{0,1\}$ indicates whether response $s_{i}$ is hallucinated ($h_{i}=1$) or not ($h_{i}=0$).
+NMISS is computed only over non-hallucinated responses.
+Formally, let
+$I=\{1,2,\dots,n\},\quad I_{\text{NH}}=\{\,i\in I\;\mid\;h_{i}=0\,\}$
+From this subset, I extract the valid cases where the traditional metric score $f_{\text{ref},i}(s_{i},r_{i})$ lies strictly between two thresholds $\tau_{\min}$ and $\tau_{\max}$, with $0\leq\tau_{\min}<\tau_{\max}\leq 1$.
+This constraint filters out trivial cases, which are perfect matches or complete mismatches:
+$I_{\text{valid}}=\{\,i\in I_{\text{NH}}\;\mid\;\tau_{\min}<f_{\text{ref},i}(s_{i},r_{i})<\tau_{\max}\,\}$
+The corresponding subset of responses is then defined as
+$\mathcal{S}_{\text{valid}}=\{\,s_{i}\;\mid\;i\in I_{\text{valid}}\,\}$
+where $f_{\text{ref},i}(s_{i},r_{i})$ denotes the score computed by a traditional reference-based metric (see Stage 1 in Algorithm 1).
+For each valid case $i\in I_{\text{valid}}$, NMISS is considered to outperform the traditional metric if
+$f_{\text{NMISS},i}(s_{i},r_{i},c_{i})>f_{\text{ref},i}(s_{i},r_{i})$
+where $f_{\text{NMISS},i}$ is the NMISS score calculated over the triplet $(s_{i},r_{i},c_{i})$, incorporating both reference alignment and contextual grounding.
+The overall percentage of NMISS outperformance across all valid cases is thus computed as
+$\frac{\big|\{\,i\in I_{\text{valid}}\;\mid\;f_{\text{NMISS},i}(s_{i},r_{i},c_{i})>f_{\text{ref},i}(s_{i},r_{i})\,\}\big|}{|I_{\text{valid}}|}\times 100$ (16)
+This formulation maintains full consistency with the operational definition of NMISS and isolates its contribution precisely where traditional metrics underperform.
+By focusing on ambiguous, semantically rich cases that lie outside trivial bounds, NMISS outperformance becomes a robust indicator of a model’s ability to generate contextually grounded yet non-literal responses.
 4 Experimental Setup
@@
 Data comprises text-based news in Italian, collected from various sources, and all focused on the health domain within Italy.
-This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues across the country.
-I collected 126,470 news articles from 2010 to 2024; their frequency is shown in Figure 1.
+This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues nationwide.
+News dataset consists of 126,470 articles from 2010 to 2024; their frequency is shown in Figure 1.
 A pre-processing step was employed to clean the raw text of HTML tags and other unwanted artifacts to ensure data quality.
@@
 Figure 1: News Frequency Aggregated by Year
-Next, I developed a list of 100 questions related to public health.
-Questions range from very generic to specific or ambiguous, leading to different levels of interpretation ease.
-The questions are categorized into three levels of understanding, based on the objective criterion of question complexity, measured by the average question length as a proxy for how detailed the question is (low: 30 characters, mid: 40 characters, high: 50 characters).
-Additionally, all questions were reviewed by a domain expert to ensure accuracy and relevance.
-All questions have been answered manually; Table 3 shows an example of questions.
+4.2 Human Evaluation Protocol
+To build a robust evaluation framework, I constructed a dataset of 100 public health questions [^1] The full list of questions is available upon request..
+The questions span a range of complexity, from broad, general topics to highly specific or ambiguous ones, yielding varying levels of interpretability.
+To systematically assess this variation, questions were grouped into three difficulty categories—low, mid, and high—based on average character length, used as a proxy for cognitive and semantic complexity [^2] Respectively, low: 30 characters; mid: 40 characters; high: 50 characters..
+Each question was reviewed by a domain expert to ensure relevance and factual accuracy, and a manual reference answer is prepared for every entry.
+Table 3 presents sample questions by category.
 Table 3: Categorized Example Questions
@@
 LOW What is a common cold?
+To evaluate hallucination detection performance, each model-generated answer was manually annotated as either hallucinated or non-hallucinated as formalized in Section 3.3.2.
+An answer was labeled as hallucinated if it contained factual inaccuracies, fabricated details, or information not supported by the retrieved context or the corresponding reference answer.
+This binary annotation served as the ground truth for computing alignment between automatic metrics and human judgment.
+The annotated dataset forms the basis for the analyses presented in Section 5, enabling direct comparison between traditional metrics and NMISS.
 4.3 User-Agent Chatbot Framework
-This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news, using a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing.
-This multi-model setup ensures robustness in handling diverse queries related to Italian health news, providing accurate, coherent, and contextually grounded responses.
-The system begins with Sentence Transformers to generate dense vector embeddings for the Document Database, which are stored and accessed via the Dense Vector Space.
-This enables efficient retrieval of relevant information through semantic search and clustering.
-After experimenting with different chunk lengths, I optimized retrieval performance by selecting a chunk length of 512 tokens with a 64-token overlap, as evaluated by human reviewers on a dataset of 100 health-related queries.
-For the retrieval part, I choose the Sentence Transformer model ‘paraphrase-multilingual-MiniLM-L12-v2‘ [^1] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for its high-quality embeddings across multiple languages, making it ideal for Italian health news [^2] After testing various chunk token sizes, the chunk length of 512 tokens with a 64-token overlap provide the best balance between retrieval performance and contextual continuity..
-Figure 2: User-Agent Application Workflow.
+This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news.
+The framework adopts a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing.
+This multi-model setup ensures robustness in managing diverse user queries, providing accurate, coherent, and contextually grounded responses.
+Figure 2: User-Agent Application Workflow
+Note.
 Text data is split into chunks for efficient retrieval, indexed using a Sentence Transformer model, and stored as dense vector embeddings in a vector store.
 The system retrieves relevant chunks based on query embeddings and passes them to an LLM for response generation.
-For the retrieval part, I choose various LLMs handling the Italian language.
-LLaMA2 and LLaMA3 are generative models developed by Meta[49, 30], optimized through supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
-Mistral[18] is a 7-billion parameter model offering competitive performance in generating accurate responses.
-GPT-4[1], the latest multimodal model from OpenAI.
-Finally, Gemma and Gemma 2[46, 47], lightweight models from Google, are fine-tuned for Italian question-answering and summarization tasks.
-The workflow illustrated in Figure 2 starts with retrieving relevant information from a large corpus of chunk news by embedding the user query into a vector space using the sentence transformer.
-The query embedding is matched against stored dense vector embeddings, and the top-K most similar document chunks are retrieved from the vector database.
-Next, the generation component uses the retrieved document chunks as context to produce a response.
-The LLM receives both the original query and the retrieved context, generating a natural language response grounded in factual data, thus reducing hallucinations.
+The system begins with a Sentence Transformer model used to generate dense vector embeddings for the document database, which are then stored and indexed within the DVS.
+This enables efficient retrieval of relevant information through semantic search and clustering.
+After experimenting with different chunking configurations, retrieval performance was optimized using a chunk length of 512 tokens with a 64-token overlap [^3] Empirically, a chunk size of 512 tokens with a 64-token overlap provided the best trade-off between retrieval accuracy and contextual continuity., evaluated on a dataset of 100 health-related queries.
+For embedding generation, I employ the Sentence Transformer model Paraphrase Multilingual MiniLM L12444https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2, selected for its high-quality multilingual embeddings, particularly suitable for Italian health news.
+For the generation component, I evaluate multiple LLMs [^5] All open-source models were evaluated with a temperature of 0.7 for comparability.
+capable of processing Italian text.
+LLaMA2 and LLaMA3, developed by Meta [touvron2023llama, meta2024llama3], are optimized through Supervised Fine-Tuning (SFT) and Reinforcement Learning with Human Feedback (RLHF).
+Mistral [jiang2023mistral] is a 7-billion-parameter model known for its efficiency and competitive generation accuracy.
+GPT-4 [achiam2023gpt] represents the latest multimodal model from OpenAI, while Gemma and Gemma2 [team2024gemma, team2024gemma2] are lightweight models from Google, fine-tuned for question-answering and summarization tasks.
+The workflow illustrated in Figure 2 begins by embedding the user query into the same vector space as the preprocessed document chunks.
+The query embedding is compared against stored vectors to retrieve the top-$K$ most semantically similar chunks from the vector database.
+These retrieved documents are then supplied to the LLM as contextual input.
+During generation, the model receives both the user query and the retrieved evidence, producing a natural language response grounded in factual data.
+This retrieval-augmented process reduces hallucinations and improves factual coherence across responses.
 4.4 Prompt Engineering Template
-I develop a question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context [^3] The list of the final set of Italian questions is available upon request..
-An English version is shown in Table 4.
+I develop a zero-shot question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context.
+The structure is shown in Table 4.
+Table 4: Prompt Engineering Structure
+Prompt Structure
+The system prompt instructs the model to behave as a helpful assistant and to generate answers exclusively based on the retrieved context.
+If the relevant information is not explicitly present in the context, the model is instructed to return a fallback message indicating insufficient information.
+Instruction Components
+(i) Define the assistant’s role, a chatbot specialized in general healthcare;
+(ii) Provide the retrieved context section containing excerpts from the database;
+(iii) Ask the model to answer the user’s question using only the provided context;
+(iv) Instruct the model to abstain from speculation or external knowledge;
+(v) Specify a fallback message for uncertain cases.
 This template ensures that the chatbot remains anchored to the information in the provided context.
-If the answer to a query is not found within the context, the chatbot will respond with: I’m sorry, I can’t help you based on the information I have”.
-Table 4: Prompt Engineering - Example (English translation)
-Prompt You are a chatbot that loves helping people!
-Given the following context section, answer the question using only the provided context.
-If you are unsure and the answer is not explicitly written in the documentation, respond by saying, ’I’m sorry, I can’t help you based on the information I have.’
-Question How can myocardial infarction be prevented?
-Context [1] The diagnosis of myocardial infarction on-site allows for rapid transportation by equipped ambulance, uncomfortable but not risky, directly to the hemodynamic lab, enabling coronary reopening at least 30 minutes earlier than if the patient arrives at the emergency room with their own means at the reference hospital.” – responsible for myocardial infarction [2] early infarction and saving lives with a simple blood test.” According to a press release from the Mauriziano’s press office.
-”In the last ten years,” the press release explains, ”the diagnosis of acute coronary syndromes has progressively shifted attention from the electrocardiogram (always and essential) to the qualitative and quantitative biochemical definition of myocardial damage.
-The new method provides a more specific, sensitive, and early definition of myocardial damage with undeniable advantages in response speed to the pathology, especially when it presents subtly and unexpectedly.
-In short, it recognizes infarction practically in real-time.” [3] ”Remembering that acute myocardial infarction (AMI) is a time-dependent pathology,” Mura concludes, ”it is essential to prevent and initiate therapy, either pharmacological or interventional in the hemodynamic room, within the timeframes established by national protocols.
-System Based on the context, there is no precise information on the symptoms indicating myocardial infarction.
-However, it mentions a new hospital technique capable of recognizing infarction in real-time, which is crucial given the time-dependent nature of the pathology.
+If the answer to a query is not found within the context, the chatbot is instructed to respond with: ”I’m sorry, I can’t help you based on the information I have”.
 5 Results
 5.1 RAG Performance Analysis
-The results are presented in Figures 3-4, showing the models’ performances across three levels of questioning.
-According to [8], a ROUGE score below 0.4 is considered low, between 0.4 and 0.5 is moderate, and above 0.5 is good.
-The Gemma model shows the weakest performance across all metrics.
-Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4, indicating significant difficulty in capturing relevant unigrams and bigrams.
-The ROUGE-L score is similarly weak, suggesting issues with structural coherence.
-Additionally, Gemma’s BLEU and METEOR scores remain low, showing poor n-gram precision and recall.
-With no EM, Gemma struggles to generate sequences that match the reference.
-In contrast, Gemma2 emerges as the top performer across all metrics.
-Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, indicating strong unigram and bigram relevance.
-The high ROUGE-L score points to solid structural coherence, while the BLEU and METEOR scores suggest excellent n-gram precision and recall.
-Gemma2 also achieves notable EM rates, ranging between 40% and 50%, highlighting its strength in generating exact matches, which is critical for tasks requiring precise language generation.
-GPT-4 delivers balanced performance across all metrics but shows variability across question levels.
-It follows closely behind Gemma2 in ROUGE, BLEU, and METEOR scores.
-However, a drop occurs at mid and high question levels, suggesting GPT-4 faces challenges with complex queries.
-Despite this, its BERTSCORE exceeds 0.9, and it also performs strongly in EM, rivaling Gemma2 in generating exact matches.
-Llama2, Llama3, and Mistral exhibit similar patterns.
-While their F1 scores for all ROUGE variants are relatively low, mainly due to weak precision, their recall scores approach those of Gemma2 and GPT-4.
-Mistral maintains consistent performance across all question levels, but the Llama models struggle more with more difficult questions.
-Overall, Gemma2 emerges as the most performant model, strictly followed by GPT-4, which remains competitive despite a slight decline in performance on complex queries.
-Mistral and the Llama models perform modestly across ROUGE, BLEU, and METEOR, with Mistral being more stable across all levels.
-Gemma lags significantly behind, particularly in ROUGE and BLEU.
-BERTSCORE remains consistently high across all models, with Gemma2 and GPT-4 achieving BERTSCOREs above 0.9.
+The results, summarized in Figures 34, present model performance across three levels of question complexity.
+Gemma demonstrates the weakest performance across all evaluation metrics.
+Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4666Following dalal2024text, ROUGE scores below 0.4 are considered low, between 0.4 and 0.5 moderate, and above 0.5 good., indicating limited ability to capture relevant unigrams and bigrams.
+The low ROUGE-L score further suggests poor structural coherence between generated and reference texts.
+BLEU and METEOR scores remain consistently low, reflecting weak lexical precision and recall, while the absence of EM confirms its difficulty in reproducing reference sequences.
 Figure 3: ROUGE Metrics by Question Levels
 Figure 4: BLEU, METEOR and EXACT MATCH Metrics by Question Levels
+In contrast, Gemma-2 achieves the strongest overall results.
+Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, demonstrating strong lexical alignment, while the high ROUGE-L value indicates superior sentence-level coherence.
+BLEU and METEOR scores reinforce this trend, highlighting consistent precision and recall advantages.
+Moreover, Gemma-2 attains the highest exact-match rate, evidencing precise adherence to reference phrasing when contextual alignment permits.
+GPT-4 follows closely behind, exhibiting balanced and reliable performance across all metrics.
+Although slight declines are observed at medium and high difficulty levels, it maintains high overall accuracy, achieving BERTScores above 0.9 and EM values comparable to Gemma-2.
+These results indicate strong factual grounding and effective adherence to the given prompts.
+Llama-2, Llama-3, and Mistral form a consistent mid-tier group.
+Their ROUGE F1 scores are generally lower, primarily due to reduced precision, though their recall remains comparable to that of the leading models.
+Among these, Mistral exhibits the most stable performance across all difficulty levels, whereas both Llama models experience sharper performance degradation as question complexity increases.
+Overall, a consistent performance hierarchy is observed across all evaluation metrics: Gemma-2 and GPT-4 lead as top performers; Llama-2, Llama-3, and Mistral form a stable middle tier; and Gemma remains the weakest model, often producing responses that lack coherence or contextual grounding.
+This ranking pattern holds across ROUGE, BLEU, BERTScore and METEOR, indicating robustness in the comparative evaluation.
+Among the individual metrics, METEOR yields consistently higher scores than BLEU for all models, as it accounts for stemming, synonymy, and paraphrase matching.
+This property allows to reward semantically equivalent expressions that BLEU tends to penalize, thereby providing a more lenient yet informative measure of linguistic adequacy.
+A similar trend is observed for BERTScore, whose values are uniformly high across models due to its semantic sensitivity.
 5.2 NMISS Performance Evaluation
-Results for BLEU and METEOR NMISS scores are shown in Tables 5-6.
-Findings report the scores against the novel metric for all the models across the question levels.
-In both cases, NMISS scores generally surpass their standard counterparts, indicating enhanced performance across different difficulty levels.
+Table 5 reports the percentage of cases in which NMISS outperforms its traditional counterparts, broken down by model and question difficulty.
+The results highlight NMISS’s added value across evaluation metrics and task complexities.
+Across all metrics, Gemma-2 benefits the most from NMISS, achieving the highest outperformance rates despite having one of the lowest hallucination frequencies.
+This suggests that its outputs, while factually grounded and semantically coherent, often diverge lexically from the reference answers.
+Conversely, GPT-4 shows smaller NMISS gains, consistent with its strong EM results reported in Section 5.1, indicating close alignment with the reference in both form and structure.
+Figures 57 further examine the relationship between hallucination rate and NMISS outperformance across BLEU, METEOR, and ROUGE variants.
+For BLEU (Figure 5), a clear positive correlation emerges: models with higher hallucination rates, such as LLaMA-2, LLaMA-3, and Mistral, also exhibit greater NMISS improvements.
+This pattern supports the view that BLEU penalizes elaborative or contextually grounded completions that deviate from reference phrasing, whereas NMISS compensates by rewarding semantic fidelity.
+Gemma-2 remains an exception, showing substantial NMISS gains despite minimal hallucinations, indicating high-quality, context-enriched responses that standard BLEU scoring undervalues.
+Table 5: NMISS Outperformance across Models and Question Levels
+Metric Model Low Mid High Overall
+BLEU gemma 0.00 0.00 0.00 0.00
+BLEU gemma2 87.50 83.33 76.47 82.43
+BLEU gpt4 50.00 28.57 40.00 39.52
+BLEU llama2 66.67 75.00 63.64 68.43
+BLEU llama3 66.67 60.00 73.33 66.67
+BLEU mistral 50.00 45.83 52.63 49.49
+METEOR gemma 0.00 0.00 0.00 0.00
+METEOR gemma2 0.00 21.43 15.00 12.14
+METEOR gpt4 0.00 0.00 0.00 0.00
+METEOR llama2 0.00 6.25 0.00 2.08
+METEOR llama3 4.76 0.00 4.35 3.04
+METEOR mistral 0.00 3.70 3.70 2.47
+ROUGE-1 F1 gemma 0.00 0.00 0.00 0.00
+ROUGE-1 F1 gemma2 81.82 85.71 80.00 82.51
+ROUGE-1 F1 gpt4 71.43 28.57 41.18 47.06
+ROUGE-1 F1 llama2 75.00 75.00 64.29 71.43
+ROUGE-1 F1 llama3 71.43 78.26 73.91 74.53
+ROUGE-1 F1 mistral 41.18 57.69 55.56 51.47
+ROUGE-2 F1 gemma 0.00 0.00 0.00 0.00
+ROUGE-2 F1 gemma2 9.09 7.69 0.00 5.59
+ROUGE-2 F1 gpt4 0.00 0.00 0.00 0.00
+ROUGE-2 F1 llama2 0.00 0.00 0.00 0.00
+ROUGE-2 F1 llama3 0.00 0.00 4.76 1.59
+ROUGE-2 F1 mistral 6.67 7.69 0.00 4.79
+ROUGE-L F1 gemma 0.00 0.00 0.00 0.00
+ROUGE-L F1 gemma2 36.36 50.00 50.00 45.45
+ROUGE-L F1 gpt4 0.00 14.29 11.76 8.68
+ROUGE-L F1 llama2 66.67 56.25 50.00 57.64
+ROUGE-L F1 llama3 52.38 43.48 56.52 50.79
+ROUGE-L F1 mistral 29.41 50.00 33.33 37.58
+  • Note.
+Scores represent the percentage of hallucination cases where NMISS outperformed the corresponding traditional metrics, expressed as the ratio of improved cases to valid hallucinations.
+Bold values highlight the highest scores per metric.
+“Overall” is the average across all questions difficulty.
+NMISS outperformance is computed over valid hallucination cases (excluding degenerate scores of 0 or 0.99).
+Figure 5: NMISS Utility vs. Hallucination Rate (BLEU)
+Figure 6: NMISS Utility vs. Hallucination Rate (METEOR)
+Figure 7: NMISS Utility vs. Hallucination Rate (ROUGE Variants)
+In the case of METEOR (Figure 6), the effect is less pronounced.
+This metric already incorporates stemming, synonymy, and paraphrase matching and naturally captures a larger share of semantic variability between system and reference outputs.
+Consequently, the set of non-matching tokens is comparatively small, reducing the influence of the contextual component in NMISS.
+Under these conditions, NMISS primarily refines borderline cases rather than producing large-scale adjustments.
+Nonetheless, Gemma-2 and LLaMA-3 still exhibit meaningful gains, suggesting that their elaborations are contextually grounded even when lexical overlap is already high.
+ROUGE analysis (Figure 7) reveals more varied patterns across its variants.
+ROUGE-1 largely mirrors BLEU’s behavior, with higher hallucination rates correlating with stronger NMISS improvements.
+ROUGE-L shows a weaker but still positive relationship, while ROUGE-2 remains mostly insensitive to NMISS adjustments, reflecting the rigidity of bigram-based evaluation in capturing nuanced or rephrased content.
+Overall, NMISS enhances evaluation in two complementary scenarios: (i) for models that hallucinate more frequently, by distinguishing spurious additions from contextually valid elaborations; and (ii) for high-performing models such as Gemma-2, by recognizing semantically grounded deviations that traditional metrics undervalue.
+These findings confirm NMISS’s complementary role and robustness across varying task difficulty levels and model qualities.
+Detailed average NMISS and baseline scores for BLEU, METEOR, and ROUGE are reported in Tables 610 in the Appendix.
+6 Conclusions
+This study investigates hallucination behavior in LLMs from a dual perspective: mitigation through a RAG architecture and evaluation refinement via the proposed NMISS.
+The experimental framework focuses on a question-answering task over Italian health news articles, simulating user–agent interactions in a sensitive knowledge domain.
+Six models—Gemma, Gemma-2, GPT-4, LLaMA-2, LLaMA-3, and Mistral—are evaluated in a zero-shot configuration with a fixed prompt template to ensure consistent generation behavior across systems.
+The analysis shows that Gemma-2 and GPT-4 achieve the strongest overall performance under traditional metrics such as BLEU, ROUGE, and METEOR.
+Gemma-2 demonstrates balanced precision and recall across all difficulty levels, making it a promising candidate for healthcare-related applications.
+GPT-4, while slightly behind on surface-based metrics, attains the highest EM accuracy, reflecting its strict prompt adherence and literal generation style.
+Mid-tier models—LLaMA-2, LLaMA-3, and Mistral—achieve comparable recall but lower precision, producing contextually relevant yet lexically diverse completions that traditional metrics tend to undervalue.
+To address this evaluative limitation, the NMISS framework is introduced to reward contextual alignment by incorporating information retrieved from the original input document.
+Focusing on non-hallucinated responses, the analysis quantifies the proportion of cases in which NMISS outperforms standard metrics, yielding several key findings.
+First, NMISS effectively identifies hidden semantic value in outputs overlooked by conventional metrics.
+Gemma-2, despite already leading in raw performance, achieves the highest NMISS outperformance across BLEU, ROUGE-1, and METEOR, demonstrating that NMISS can recognize semantically rich yet lexically divergent answers that remain contextually faithful.
+Second, although GPT-4 performs strongly overall, its NMISS improvement is modest.
+This reflects its rigid zero-shot generation behavior, which closely mirrors the prompt and minimizes lexical variation, leaving less room for NMISS to detect additional contextual contributions.
+Conversely, mid-tier models exhibit substantial NMISS gains, particularly for BLEU and ROUGE-1.
+Their responses, while more syntactically variable, frequently contain relevant contextual cues that overlap-based metrics penalize but NMISS rewards.
+This suggests that NMISS captures a compensatory strategy: when these models struggle to respond directly, they often enrich their outputs with contextually grounded elaborations that enhance informativeness.
+A positive correlation emerges between hallucination rate and NMISS outperformance for BLEU and ROUGE-1, reinforcing the idea that NMISS excels in borderline cases where literal matching fails to capture valid, context-driven reasoning.
+Notably, Gemma-2 remains an exception, showing that even factually accurate responses can benefit from a more context-sensitive evaluation.
+Overall, the findings demonstrate that while RAG architectures effectively mitigate hallucinations by grounding LLM outputs, NMISS complements this process by refining how output quality is assessed.
+It compensates for the underestimation introduced by traditional overlap-based metrics, which often assign lower scores to responses that are semantically correct but lexically divergent.
+By rewarding contextually aligned and informative outputs, NMISS provides a fairer evaluation of model performance.
+This dual contribution is particularly relevant for high-stakes domains such as healthcare, where contextual accuracy and informativeness often outweigh surface-level similarity.
+In conclusion, NMISS represents a valuable addition to the LLM evaluation toolkit.
+It enhances fairness and interpretability in performance assessment, particularly for models producing rich and context-aware responses.
+While top-performing models benefit from NMISS refinement, mid-tier models gain even more, as the metric recovers value otherwise overlooked by standard approaches.
+Future work could extend NMISS to multilingual and multi-hop question answering, adaptive fine-tuning pipelines, and reinforcement learning frameworks that align model outputs with context-based quality objectives.
+Beyond the Italian healthcare domain, the combined RAG–NMISS methodology generalizes to other languages and knowledge domains where factual grounding and contextual relevance are essential.
+7 Appendix
 Table 6: BLEU versus BLEU NMISS scores
@@
 gemma2 0.654 0.578 0.604 0.668 0.626 0.664
-33footnotetext: Note: Average scores. Higher values are highlighted in bold.
-For both metrics, models like GPT-4 show minimal differences between standard and NMISS scores, indicating high consistency in performance across both metrics.
-However, models like Llama2, Llama3 and Mistral exhibit more significant improvements in their BLEU NMISS scores, suggesting that NMISS better captures the models’ ability to provide additional contextual information, especially in more challenging tasks.
-Notably, Gemma2 shows strong performance overall, with NMISS scores slightly higher than standard ones, particularly at mid and high levels, highlighting its ability to maintain accuracy and detail in complex scenarios.
-The increase in NMISS scores when models struggle to answer could be attributed to how NMISS rewards the inclusion of additional relevant information or contextual details.
-When a model does not provide a concise answer, it often compensates by providing surrounding context, explanations, or related information.
-While not a perfect match to the expected answer, this additional content contributes meaningfully to the overall response.
+  • Note: Average scores. Higher values are highlighted in bold.
 Table 7: METEOR versus METEOR NMISS scores
@@
 gemma2 0.718 0.675 0.724 0.723 0.682 0.731
-33footnotetext: Note: Average scores. Higher values are highlighted in bold.
-Tables 7-9 compares ROUGE and ROUGE NMISS scores across different models and difficulty levels.
-Similar to the trends observed with BLEU and METEOR, the ROUGE NMISS scores generally outperform the standard ROUGE metrics, especially at higher difficulty levels.
-This consistent pattern across different metrics suggests that NMISS captures additional relevant information, rewarding models for providing more detailed and contextually accurate responses.
-Specifically, for most models, ROUGE NMISS scores surpass the standard ROUGE-1 and ROUGE-L scores across all difficulty levels.
-This is consistent with the behavior observed for BLEU and METEOR.
-Interestingly, the ROUGE-2 NMISS scores do not consistently outperform the standard ROUGE-2 scores across all models and levels.
-This deviation from the trend seen in ROUGE-1 and ROUGE-L might be attributed to the focus of ROUGE-2 on bigram overlaps, which are less sensitive to the additional contextual information.
+  • Note: Average scores. Higher values are highlighted in bold.
 Table 8: ROUGE-1 versus ROUGE-1 NMISS scores
@@
 HIGH gemma2 0.706 0.755 0.715 0.880 0.758 0.790
+  • Note: Average scores. Higher values are highlighted in bold.
 Table 9: ROUGE-2 versus ROUGE-2 NMISS scores
@@
 HIGH gemma2 0.635 0.664 0.641 0.639 0.664 0.642
+  • Note: Average scores. Higher values are highlighted in bold.
 Table 10: ROUGE-L versus ROUGE-L NMISS scores
@@
 HIGH gemma2 0.677 0.718 0.684 0.707 0.718 0.698
-To provide a broad perspective on the various scores, Figures 5-6 illustrate the percentage of instances where NMISS outperforms standard metrics for all questions.
-For ROUGE-1 and ROUGE-L, NMISS consistently performs better, reflecting its ability to capture more contextually rich and detailed responses.
-This further highlights NMISS’s effectiveness in evaluating model performance, particularly in greater complexity scenarios.
-The performance of ROUGE NMISS for precision across all question levels is as follows: for GPT-4 ROUGE NMISS outperforms ROUGE-1 by approximately 26% at the low level, 30% at the mid level, and 40% at the high level.
-ROUGE-L shows a similar improvement at all levels, while ROUGE-2 exhibits no notable change.
-For Llama2 and Llama3, ROUGE NMISS shows remarkable improvement in ROUGE-1, with gains between 80% and 85% across both low and high levels.
-ROUGE-L demonstrates more modest increases, ranging from 40% to 60%.
-For Mistral, NMISS provides more modest improvements, ranging from 4% to 65% in ROUGE-1 and similar results for ROUGE-L.
-Gemma2 follows a similar pattern to GPT-4, with modest improvements across all levels.
-At the same time, Gemma shows little to no improvement with NMISS, indicating that its issues may stem from machine-related problems rather than contextual relevance.
-For recall, similar trends are observed, though with different magnitudes.
-This is expected, as recall is more sensitive to the length of the context, but they remain noteworthy.
-GPT-4 gains 18% for ROUGE-1 at mid-level questions, Llama3 improves by 11% at low and 12% at high levels, and Mistral improves by 6% at the low level.
-Gemma shows no change at the low level but gains 6.1% at mid and high levels.
-The same patterns for the models across the question levels apply to METEOR and BLEU percentage analysis, where the difference stands in magnitude.
-While NMISS heavily impacts BLEU, METEOR is less pronounced.
-Figure 5: Percentage of Outperforming of ROUGE NMISS scores
-Figure 6: Percentage of Outperforming of BLEU AND METEOR NMISS scores
-Overall, the NMISS algorithm demonstrates the greatest improvement at the highest difficulty level, where the gap between NMISS and traditional metrics like BLEU and METEOR is most pronounced.
-This suggests that NMISS is particularly effective in advanced scenarios that require deeper comprehension and more complex information retrieval or when the model’s predictive power is limited.
-Notably, with GPT-4 or Gemma2 being the top performers, they benefit the least from NMISS, indicating that answers are already highly contextual and do not need additional elaboration.
-In contrast, middle-range models like Llama2, Llama3, and Mistral show the most significant gains, as NMISS helps to capture their ability to provide richer contextual information.
-Gemma sees little to no improvement, indicating that its issues are likely machine-related rather than due to contextual relevance.
-These results suggest that when models struggle to provide exact answers, they compensate by offering more detailed contextual information.
-Importantly, this does not result in hallucinations but rather reflects the models’ ability to maintain accuracy and completeness in the surrounding context.
-In such cases, even peripheral contexts can contribute to higher NMISS scores, improving the overall response quality.
-Conversely, where questions are more straightforward, the gap between NMISS and traditional metrics narrows.
-This indicates that less contextual elaboration is required, and the models perform similarly across both metrics.
-In summary, NMISS tends to reflect the models’ ability to provide richer, more informative answers in complex situations, which is less evident at lower difficulty levels where responses tend to be more concise and direct, reducing the gap between NMISS and traditional metrics.
-6 Conclusions
-In this paper, I investigate hallucinations in LLMs using a combination of detection and mitigation techniques.
-I evaluate a user-agent chatbot interaction with Italian health news articles through a question-answering framework based on the RAG architecture.
-Models including Gemma, Gemma2 GPT-4, Llama2, Llama3, and Mistral are analyzed.
-To enhance traditional evaluation metrics like ROUGE, METEOR, and BLEU, I introduce the NMISS scoring system, which emphasizes contextual relevance in model-generated responses.
-NMISS allows the identification of instances where traditional metrics undervalue contextually rich answers that deviate from the reference.
-Regarding performance, Gemma2 and GPT-4 consistently outperformed other models.
-Gemma2 demonstrated strong results across ROUGE, METEOR, and BLEU, making it a top choice for healthcare-related chatbot applications.
-GPT-4 excelled in EM accuracy, highlighting its ability to generate sequences closely aligned with reference answers.
-The mid-range models Llama2, Llama3, and Mistral show similar performance patterns.
-Although their F1 scores across all ROUGE variants are lower due to weaker precision, their recall scores are close to those of Gemma2 and GPT-4, indicating their ability to capture a significant amount of relevant content despite generating less precise outputs.
-Mistral maintains consistent performance across all difficulty levels, while Llama2 and Llama3 struggled with more complex queries.
-NMISS application reveals that when a model struggles to provide exact answers, it often compensates by delivering richer contextual information.
-This does not result in hallucinations but highlights the LLM’s capacity to provide valuable information embedded in the context, even when direct answers are challenging to generate.
-This demonstrates the importance of context as a carrier of valuable information.
-NMISS consistently outperforms traditional ROUGE-1 and ROUGE-L scores across all question levels, and similar trends are observed for BLEU and METEOR.
-It is important to note that high-performing models like Gemma2 and GPT-4 benefit less from NMISS, which is expected.
-These models already provide highly precise and contextually accurate responses.
-In contrast, NMISS proves particularly beneficial for mid-tier models like Llama2, Llama3, and Mistral, which often generate contextually richer responses when faced with more challenging queries.
-These models benefit from NMISS because it captures and rewards their ability to provide additional relevant context, even if their precision is lower.
-On the other hand, Gemma shows slight improvement across all metrics, suggesting that its hallucination issues are machine-intrinsic and less related to context-awareness, which NMISS is designed to evaluate.
-While RAG models mitigate hallucinations by grounding outputs in external data, NMISS refines the evaluation process.
-It helps distinguish “false positives” — cases where traditional metrics flag hallucinations due to deviations from the reference, even though the response is contextually accurate.
-NMISS reduces penalties for models that provide relevant contextual information, even when their answers do not match the reference.
-This distinction underlines the complementary nature of RAG and NMISS: RAG acts as a tool for preventing hallucinations, while NMISS improves how I evaluate context and relevance in model output, serving as a diagnostic tool for better understanding hallucinations.
-Future research could explore extending this framework to more complex datasets, tasks, and languages, further advancing hallucination mitigation and improving LLM performance in real-world applications.
-Acknowledgements
-This publication was produced while attending the PhD program in Economics and Business at the University of Cagliari, Cycle 38, with the support of a scholarship financed by the Ministerial Decree no.
-351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”.
+  • Note: Average scores. Higher values are highlighted in bold.